### v3

# Exponentiations of ultrafilters

Abstract
In recent years, several problems regarding the partition regularity of exponential configurations have been studied in the literature, in some cases using the properties of specific ultrafilters. In this paper, we start to lay down the foundations of a general theory of exponentiations of ultrafilters, with a particular focus on their combinatorial properties and the existence of idempotents.
Mathematics subject classification 2020: 05D10, 22A15, 54D35.
Keywords: Stone-Čech compactification, idempotents, exponential triples, partition regularity.
1   Introduction
A large class of problems in arithemtic Ramsey theory regards the so-called partition regular configurations and equations on =$\mathbb{N}=\{1,2,\dots\}$.
Definition 1.1.
Let $f_{1}\left(x_{1},\dots,x_{n}\right),\dots,f_{m}\left(x_{1},\dots,x_{n}\right):\mathbb{N}^{n}\rightarrow\mathbb{N}$. We say that the configuration $\{f_{1}\left(x_{1},\dots,x_{n}\right),\dots,f_{m}\left(x_{1},\dots,x_{n}\right)\}$ is partition regular on $\mathbb{N}$ (PR from now on) if for all finite partitions $\mathbb{N}=A_{1}\cup\dots\cup A_{k}$ there are $i\leq k$ and $x_{1},\dots,x_{n}\in\mathbb{N}$ such that for all $j\leq m$ $f_{j}\left(x_{1},\dots,x_{n}\right)\in A_{i}$.
Analogously, we say that the system $\sigma\left(x_{1},\dots,x_{n}\right)$ defined by
$\begin{cases}f_{1}\left(x_{1},\dots,x_{n}\right)=0,\\ \ \ \ \ \ \ \ \ \ \vdots\\ f_{m}\left(x_{1},\dots,x_{n}\right)=0\end{cases}$
is PR if for all finite partitions $\mathbb{N}=A_{1}\cup\dots\cup A_{k}$ there are $i\leq k$ and $x_{1},\dots,x_{n}\in A_{i}$ such that $\sigma\left(x_{1},\dots,x_{n}\right)=0$.
In [7], Schur proved one of the first and most important results in Arithmetic Ramsey Theory.
Theorem 1.2 (Additive Schur Theorem).
The configuration $\{x,y,x+y\}$ is PR.
Equivalently, the additive Schur Theorem states that the equation $x+y=z$ is PR.
Applying transformations to partitions allows to easily deduce new results by just considering the shape of transformed partition regular configurations. A useful trick, that sometimes allows turning additive structures in multiplicative structures, and multiplicative structures in exponential ones, is that of considering $\log_{2}$-base partitions: for $B\subseteq\mathbb{N}$, let $2^{-B}=\{n\in\mathbb{N}\mid 2^{n}\in B\}$ and, given a partition $\mathbb{N}=A_{1}\cup\dots\cup A_{k}$, let the associated $\log_{2}$-base partition be given by $\mathbb{N}=2^{-A_{1}}\cup\dots\cup 2^{-A_{k}}$. For example, given any finite partition of $\mathbb{N}$, by Theorem 1.2 in the $\log_{2}$-base partition we find monochromatic $x,y,z$ with $x+y=z$; hence, $2^{x},2^{y},2^{z}=2^{x+y}=2^{x}\cdot 2^{y}$ were monochromatic with respect to the original partition. This proves the multiplicative version of Schur.
Theorem 1.3 (Multiplicative Schur Theorem).
The configuration $\{x,y,x\cdot y\}$ is PR.
Equivalently, the multiplicative Schur Theorem states that the equation $xy=z$ is PR.
Recently, in [8], the question about the partition regularity of $\{x,y,x^{y}\}$, which we will call an exponential triple, was posed and partially answered, by using ultrafilters, by Sisto; its full solution was first given in [6] by Sahasrabudhe who, among other generalizations, proved the exponential Schur Theorem [^2] At the best of my knowledge, this theorem has not been called explicitly like this in the literature before.:
Theorem 1.4 (Exponential Schur Theorem).
The configuration $\{x,y,x^{y}\}$ is PR.
Again, the above theorem can be rephrased in terms of equations: it states that $x^{y}=z$ is PR. Sahasrabudhe’s proof used repeated applications of van der Waerden theorem and the compactness property on the convergence of subsequences in the space of $k$-partitions of $\mathbb{N}$. One of the key ingredients in his proof was the use of the function $f(n,m):=2^{n}m$, combined with associated $\log_{2}$-base partitions. Starting from observations about this function, in [2, Theorem 1.1] Di Nasso and Ragosta found an ultrafilters proof of the partition regularity of the configuration $\{x,y,x2^{y}\}$, from which the exponential Schur’s theorem can be immediately deduced by the same $\log_{2}$-partition trick that allows proving the multiplicative version of Schur Theorem from the additive one (see [2, Corollary 1.3]). Their proof was based on a clever trick based on tensor products of ultrafilters [^3] Through the paper we assume the reader to have a basic knowledge of the main properties of $\beta\mathbb{N}$; we refer to the monograph [5] for an expensive presentation of ultrafilters and their applications to combinatorics. and the extension to $\left(\beta\mathbb{N}\right)^{2}$ of the same function $f(n,m)=2^{n}m$ used by Sahasrabudhe is applied. Di Nasso and Ragosta used ultrafilters methods also to study the partition regularity of infinite exponential configurations in [3]. As its often the case with partition regularity problems, the ultrafilters approach provides a simple and informative proof that allows for several generalizations: this is a common situation, as proven by an enormous literature on the subject (instead of citing here hundreds of papers, we refer to [5] and its bibliography).
Our main goal is to start a systematic study of which ultrafilters are related to exponential triples, and what other combinatorial properties such ultrafilters have. This will be done by introducing two different exponentiations of ultrafilters, and by studying their properties. In particular, we will prove that all multiplicatively piecewise syndetic sets contain exponential triples which, combined with our exponentiations of ultrafilters, will allow to show that large families of ultrafilters witness the exponential Schur Theorem, in the sense that all their sets contain exponential triples.
When a binary operations on $\beta\mathbb{N}$ is introduced, it is natural to search for idempotents since, as it is well-known since Galvin and Glazer’s proof of Hindman Theorem, such idempotents have plenty of interesting combinatorial properties. Henceforth, we study and solve negatively (completely for one extension, partially for the other) the problem of the existence of idempotents in Section 2. It turns out that idempotents for such exponentiations are strictly related to idempotents for other exponential-like operations, namely generalizations of the operation $(p,q)\rightarrow 2^{p}\odot q$, which first appeared in [2].
The novelty of most of our results generates a large number of open questions and future possible directions. We conclude the paper with a discussion of many of these in Section 3.
2   Exponentiations of ultrafilters
As ultrafilters are the key notion of this paper, we assume the reader to have a basic knowledge of the algebra of $\beta\mathbb{N}$. We just briefly recall a few definitions, mostly so that we can fix a few notations.
A notion that we will use often is that of tensor product of ultrafilters in $\beta\mathbb{N}$, defined as follows:
Definition 2.1.
Given $p,q\in\beta\mathbb{N}$, the tensor product $p\otimes q$ of $p,q$ is the ultrafilter on $\beta(\mathbb{N}^{2})$ defined as follows: for all $A\subseteq\mathbb{N}^{2}$
$A\in p\otimes q\Leftrightarrow\{a\in\mathbb{N}\mid\{b\in\mathbb{N}\mid(a,b)\in A\}\in q\}\in p.$
Moreover, we will need to talk about extensions of functions to $\beta\mathbb{N}$: for $f:\mathbb{N}^{n}\rightarrow\mathbb{N}$, we will denote by $\widehat{f}$ its unique continuous extension $\widehat{f}:\beta(\mathbb{N}^{n})\rightarrow\beta\mathbb{N}$. As usual, to simplify notations we identify each natural number $n\in\mathbb{N}$ with the principal idempotent it generates, so that $\mathbb{N}\subseteq\beta\mathbb{N}$.
The only purely combinatorial notion that we will need is that of multiplicatively piecewise set, defined as follows. A set $A\subseteq\mathbb{N}$ is multiplicatively syndetic if there exists a finite set $F$ such that for all $n\in\mathbb{N}$ there exists $f\in F$ such that $fx\in A$; $A$ is multiplicatively thick if for all $F$ finite subset of $\mathbb{N}$ there exists $x\in\mathbb{N}$ such that $xF\subseteq A$; $A$ is multiplicatively piecewise syndetic if $A=B\cap C$ where $B$ is multiplicatively syndetic and $C$ is multiplicatively thick. The relevant property of multiplicatively piecewise sets for us is the following:
Proposition 2.2.
[5, Theorem 4.40] $A\subseteq\mathbb{N}$ is multiplicatively piecewise syndetic if and only if there exists $p\in K(\beta\mathbb{N},\odot)$ such that $A\in p$.
Two extensions of exponentiation to ultrafilters
As it is well known, there are two natural ways to extend the sum from $\mathbb{N}$ to $\beta\mathbb{N}$: the usual one $(p,q)\rightarrow p\oplus q$ defined by letting, for every $A\in\mathbb{N}$,
$A\in p\oplus q\Leftrightarrow\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid n+m\in A\}\in q\}\in p,$
and the reverse sum $p\tilde{\oplus}q:=q\oplus p$; if $S(n,m):=n+m$ denotes the sum of natural numbers, then $p\oplus q=\widehat{S}(p\otimes q)$ and $q\oplus p=\widehat{S}(p\otimes q)$. Whilst different, $\oplus$ and $\tilde{\oplus}$ have the same properties, modulo properly translating their meaning by switching “left” with “right” notions. This is a consequence of the sum being commutative on $\mathbb{N}$.
Exponentiation, on the other hand, is not commutative nor associative. Hence, when extending it to $\beta\mathbb{N}$, we have at least two natural and truly distinct possible definitions to consider [^4] Not including those that could be obtained trivially by reversing the order of the variables in the operations..
Definition 2.3.
Given $p,q\in\beta\mathbb{N}$ we let
$E_{1}(p,q):=\{A\subseteq\mathbb{N}\mid\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid n^{m}\in q\}\in p\},$
and
$E_{2}(p,q):=\{A\subseteq\mathbb{N}\mid\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid m^{n}\in q\}\in p\}.$
Equivalently, if $e:\mathbb{N}^{2}\rightarrow\mathbb{N}$ is the exponentiation $e(n,m)=n^{m}$, then $E_{1}(p,q)=\widehat{e}(p\otimes q)$, whilst $E_{2}(p,q)=\widehat{e}(q\otimes p)$.
For $n,m\in\mathbb{N}$, trivially we have that $E_{1}(n,m)=E_{2}(m,n)=n^{m}$; more in general, for $n\in\mathbb{N},p\in\beta\mathbb{N}$ we have that $E_{1}(n,p)=E_{2}(p,n):=p^{n}$ and $E_{1}(p,n)=E_{2}(n,p)=n^{p}$ as, for all $A\subseteq\mathbb{N}$
$A\in E_{1}(n,p)\Leftrightarrow\{m\in\mathbb{N}\mid n^{m}\in A\}\in p\Leftrightarrow A\in E_{2}(p,n),$
$A\in E_{1}(p,n)\Leftrightarrow\{m\in\mathbb{N}\mid m^{n}\in A\}\in p\Leftrightarrow A\in E_{2}(n,p).$
Again, in terms of extension of functions, the above shows that, for $n$ fixed, $p^{n}=\widehat{e_{n}}(p)$, where $e_{n}:\mathbb{N}\rightarrow\mathbb{N}$ is the function that maps $m$ to $m^{n}$, and $n^{p}=\widehat{e^{\prime}_{n}}(p)$, where $e^{\prime}_{n}:\mathbb{N}\rightarrow\mathbb{N}$ is the function that maps $m$ to $n^{m}$.
Notation: we will continue to use the simplified notation $n^{p}$ to denote any of $E_{1}(n,p),E_{2}(p,n)$, and we will use the simplified notation $p^{n}$ to denote any of $E_{1}(p,n),E_{2}(n,p)$.
At first sight, it might look like the equality $E_{1}(p,q)=E_{2}(q,p)$ could be true for all $p,q\in\beta\mathbb{N}$; however, this is not the case. In fact, for example, let both $p,q$ contain the set $P$ of primes. Then
$P_{1}=\{a^{b}\mid a,b\in P,a<b\}\in E_{1}(p,q),P_{2}=\{a^{b}\mid a,b\in P,a>b\}\in E_{2}(p,q)$
but $P_{1}\cap P_{2}=\emptyset$, so $E_{1}(p,q)\neq E_{2}(q,p)$. We will provide some examples of nonprincipal ultrafilters $p,q$ satisfying $E_{1}(p,q)=E_{2}(q,p)$ after Proposition 2.4.
$E_{1}$ and $E_{2}$ allow to extend, in a proper way, several algebraical properties of exponentiation from natural numbers to ultrafilters. We list here those that we will need in the rest of the paper. To avoid cumbersome reformulations, some of the properties proven in the following proposition are stated for rational numbers and ultrafilters on the rationals, to which the exponentiation functions $E_{1},E_{2}$ can be extended trivially.
Proposition 2.4.
Let $p,q,r\in\beta\mathbb{N}\setminus\{0,1\},n,m\in\mathbb{N}\setminus\{1\},a\in\mathbb{Q}\setminus\{0\}$. The following properties hold:
  1. 1.
    $E_{1}\left(p^{a},q\right)=E_{1}\left(p,aq\right)$ and $E_{2}\left(q,p^{a}\right)=E_{2}\left(aq,p\right)$;
  2. 2.
    $n^{p}=n^{q}\Leftrightarrow p=q$ and $p^{n}=q^{n}\Leftrightarrow p=q$;
  3. 3.
    $n^{p}=m^{p}\Leftrightarrow n=m$ and $p^{n}=p^{m}\Leftrightarrow n=m$;
  4. 4.
    $n^{p}\odot n^{q}=n^{p\oplus q}$, but $E_{1}(r,p)\odot E_{1}(r,q)\neq E_{1}(r,p\oplus q)$ and $E_{2}(p,r)\odot E_{2}(q,r)\neq E_{2}(p\oplus q,r)$ in general;
  5. 5.
    $E_{1}\left(p,q\odot r\right)=E_{1}\left(E_{1}(p,q),r\right)$; in particular, if $q$ is multiplicatively idempotent, $E_{1}(p,q)=E_{1}\left(E_{1}(p,q),q\right)$;
  6. 6.
    $E_{2}\left(p,E_{2}(q,r)\right)=E_{2}\left(p\odot q,r\right)$; in particular, if $p$ is multiplicatively idempotent, $E_{2}\left(p,E_{2}(p,q)\right)=E_{2}(p,q)$.
Proof.
All these proofs consist, basically, of straightforward checks on the definitions.
$(1)$ Let $A\subseteq\mathbb{N}$. Then
$A\in E_{1}\left(p^{a},q\right)\Leftrightarrow\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid n^{m}\in A\}\in q\}\in p^{a}\Leftrightarrow$
$\{n\in\mathbb{N}\mid n^{a}\in\{l\in\mathbb{N}\mid\{m\in\mathbb{N}\mid l^{m}\in A\}\in q\}\}\in p\Leftrightarrow\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid\left(n^{a}\right)^{m}\in A\}\in q\}\in p\Leftrightarrow$
$\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid\left(n^{am}\right)\in A\}\in q\}\in p\Leftrightarrow\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid\left(n^{m}\right)\in A\}\in aq\}\in p\Leftrightarrow$
$A\in E_{1}\left(p,aq\right).$
The $E_{2}$-case is analogous.
$(2)$ If $p=q$, both equivalences are trivial.
If $n^{p}=n^{q}$ it means that for all $A\subseteq\mathbb{N}$, if $B=\{a\in\mathbb{N}\mid n^{a}\in A\}$ then $B\in p\Leftrightarrow B\in q$, which shows that if we let $P_{n}=\{a\in\mathbb{N}\mid\exists b\in\mathbb{N}\ a=b^{n}\}$ and $log_{n}:P_{n}\rightarrow\mathbb{N}$ be the $n$-base logarithm and $\widehat{log_{n}}$ be its extension to ultrafilters containing $P_{n}$ then $p=\widehat{log_{n}}\left(n^{p}\right)=\widehat{log_{n}}\left(n^{q}\right)=q$.
The other equivalence can be proven similarly.
$(3)$ If $n=m$, both equivalences are trivial.
If $n^{p}=m^{p}$, first we observe that $n^{\mathbb{N}}=\{n^{k}\mid k\in\mathbb{N}\}\in n^{p}$ and $m^{\mathbb{N}}=\{m^{k}\mid k\in\mathbb{N}\}\in m^{p}$. Therefore, if $n^{p}=m^{p}$ we have that $n^{\mathbb{N}}\cap m^{\mathbb{N}}\in n^{p}$. Hence $n^{\mathbb{N}}\cap m^{\mathbb{N}}\neq\emptyset$, which forces $n^{a}=m$ for some $a\in\mathbb{Q}$, so $n^{p}=n^{ap}$ from $(1)$, which from $(2)$ forces $p=ap$, which forces $a=1$, so $n=m$.
The other equivalence is similar.
$(4)$ First, let us show that $n^{p}\odot n^{q}=n^{p\oplus q}$: let $A\subseteq\mathbb{N}$. Then
$A\in n^{p}\odot n^{q}\Leftrightarrow\{a\in\mathbb{N}\mid\{b\in\mathbb{N}\mid ab\in A\}\in n^{q}\}\in n^{p}\Leftrightarrow$
$\{d\in\mathbb{N}\mid n^{d}\in\{a\in\mathbb{N}\mid\{c\in\mathbb{N}\mid n^{c}\in\{b\in\mathbb{N}\mid ab\in A\}\in q\}\}\}\in p\Leftrightarrow$
$\{d\in\mathbb{N}\mid\{c\in\mathbb{N}\mid n^{c}\cdot n^{d}\in A\}\in q\}\in p\Leftrightarrow\{d\in\mathbb{N}\mid\{c\in\mathbb{N}\mid n^{c+d}\in A\}\in q\}\in p\Leftrightarrow$
$\{e\in\mathbb{N}\mid n^{e}\in A\}\in p\oplus q\Leftrightarrow A\in n^{p\oplus q}.$
For the other claims: let $P=\{n\in\mathbb{N}\mid n\ \text{is a prime}\}$ and let $P\in r$. Let $p,q$ be nonprincipal. Then a direct computation shows that
$I_{1}=\{n\in\mathbb{N}\mid\exists p_{1},p_{2}\in P,a,b\in\mathbb{N}\ \text{s.t.}\ p_{1}<a<p_{2}<b\ \text{and}\ p_{1}^{a}p_{2}^{b}=n\}\in E_{1}(r,p)\odot E_{1}(r,q),$
$I_{2}=\{n\in\mathbb{N}\mid\exists p_{1},p_{2}\in P,a,b\in\mathbb{N}\ \text{s.t.}\ p_{1}>a>p_{2}>b\ \text{and}\ p_{1}^{a}p_{2}^{b}=n\}\in E_{2}(p,r)\odot E_{2}(q,r),$
whilst
$J=\{n\in\mathbb{N}\mid\exists p\in P,a\in\mathbb{N}\ \text{s.t.}\ p^{a}=n\}\in E_{1}(r,p\oplus q)\cap E_{2}(p\oplus q,r),$
and we conclude as $I_{1}\cap J=I_{2}\cap J=\emptyset.$
$(5)$ Let $A\subseteq\mathbb{N}$. Then
$A\in E_{1}(p,q\odot r)\Leftrightarrow\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid n^{m}\in A\}\in q\odot r\}\in p\Leftrightarrow$
$\{n\in\mathbb{N}\mid\{a\in\mathbb{N}\mid\{b\in\mathbb{N}\mid n^{ab}\in A\}\in r\}\in q\}\in p\Leftrightarrow\{n\in\mathbb{N}\mid\{a\in\mathbb{N}\mid\{b\in\mathbb{N}\mid\left(n^{a}\right)^{b}\in A\}\in r\}\in q\}\in p\Leftrightarrow$
$\{n\in\mathbb{N}\mid\{a\in\mathbb{N}\mid\{b\in\mathbb{N}\mid\left(n^{a}\right)^{b}\in A\}\in r\}\in q\}\in p\Leftrightarrow\{c\in\mathbb{N}\mid\{d\in\mathbb{N}\mid c^{d}\in A\}\in r\}\in E_{1}\left(p,q\right)\Leftrightarrow$
$A\in E_{1}\left(E_{1}\left(p,q\right),r\right).$
$(6)$ This is analogous to $(5)$. Let $A\subseteq\mathbb{N}$. Then
$A\in E_{2}\left(p,E_{2}(q,r)\right)\Leftrightarrow\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid m^{n}\in A\}\in E_{2}(q,r)\}\in p\Leftrightarrow$
$\{n\in\mathbb{N}\mid\{a\in\mathbb{N}\mid\{b\in\mathbb{N}\mid\left(b^{a}\right)^{n}\in A\}\in r\}\in q\}\in p\Leftrightarrow\{n\in\mathbb{N}\mid\{a\in\mathbb{N}\mid\{b\in\mathbb{N}\mid b^{an}\in A\}\in r\}\in q\}\in p\Leftrightarrow$
$\{c\in\mathbb{N}\mid\{b\in\mathbb{N}\mid b^{c}\in A\}\in r\}\in p\odot q.\qed$
Example 2.5.
Let $n\in\mathbb{N},n>1$ and $p,q\in\beta\mathbb{N}\setminus\{0\}$ be such that $p\odot q=q\odot p$. Then by Proposition 2.4.we have that
$E_{1}\left(n^{p},q\right)=n^{p\odot q}=n^{q\odot p}=E_{2}\left(q,n^{p}\right).$
In particular, this shows the existence of infinitely many pairs of nonprincipal ultrafilters solving $E_{1}(p,q)=E_{2}(q,p)$ (just take, e.g., $p=2^{q}$).
Their definition should make evident that $E_{1},E_{2}$ are ultrafilters operations with relationships with exponential configurations.
Example 2.6.
Let $p\in\beta\mathbb{N}\setminus\mathbb{N}$ and let $q=E_{1}(p,p)$. Let $A\in q$. Then
$B_{0}=\{m\in\mathbb{N}\mid\{n\in\mathbb{N}\mid\{m^{n}\in A\}\in p\}\in p.$
Now take $F_{1}\in\mathcal{P}_{f}(B_{0})$ and let
$B_{1}=\bigcap_{t\in F_{1}}\{n\in\mathbb{N}\mid t^{n}\in A\}\cap B_{0}\in p.$
Proceeding inductively like above, we construct a sequence $\langle F_{i}\rangle_{i=1}^{\infty}$ of finite subsets of $B_{0}$ with the property that for all $j<i$ one has that
$F_{j}^{F_{i}}=\{m^{n}\mid m\in F_{j},n\in F_{i}\}\in A.$
For example, if $F_{i}=\{a_{i}\}$ for all $i$, we obtain an infinite subset $X=\{a_{i}\mid i\in\mathbb{N}\}$ of $B_{0}$ such that $\{a_{j}^{a_{i}}\mid a_{i},a_{j}\in X,j<i\}\subseteq A$.
Moreover, by choosing $p$ with additional properties, one could get more complicated configurations. For example, by taking $p$ such that each set in $p$ contains arbitrarily long arithmetic progressions, we can take all sets $F_{i}$ to be arithmetic progressions of length $i$ in $B$.
Example 2.7.
If we argue as above with $E_{2}$ instead of $E_{1}$, we get analogous results with the order relationship between $i$ and $j$ switched. So, in the same notations, we get a sequence $\langle F_{i}\rangle_{i=1}^{\infty}$ of finite subsets of $B_{0}$ with the property that for all $j<i$ one has that
$F_{j}^{F_{i}}=\{m^{n}\mid m\in F_{j},n\in F_{i}\}\in A.$
For example, if $F_{i}=\{a_{i}\}$ for all $i$, we obtain an infinite subset $X=\{a_{i}\mid i\in\mathbb{N}\}$ of $B_{0}$ such that $\{a_{j}^{a_{i}}\mid a_{i},a_{j}\in X,j>i\}\subseteq A$.
The reader with a basic knowledge of ultrafilters will recognize that a trivial way to strengthen the above examples would be to take $p$ such that $E_{1}(p,p)=p$ or $E_{2}(p,p)=p$, namely some kind of exponential idempotent. We devote Section 2.2 to the study of the existence of such ultrafilters.
To conclude this section, we want to show that the operations $E_{1},E_{2}$ are related to exponential triples. This should not come as a surprise as, even if using a different language, Di Nasso and Ragosta in [2] to prove the exponential Schur theorem employed ultrafilters $u$ of the form $u=2^{q}$ where $q$ had the form $q=2^{p}\odot p$ for some $p$ witness of van der Waerden Theorem. In [4], the authors proved the following result, which allows to generalize the observation of Di Nasso and Ragosta::
Theorem 2.8.
Let $A$ be a multiplicatively piecewise syndetic set. Then $A$ contains an exponential triple.
As an immediate corollary, the following holds.
Corollary 2.9.
Let $\mathcal{U}\in\overline{K(\beta\mathbb{N},\odot)}$. Then every $A\in\mathcal{U}$ contains an exponential triple.
Notice that, by Corollary 2.9, it is simple to prove that several configurations are partition regular by using [1, Lemma 2.1]. To recall it, we need to introduce the following notation: for $p\in\beta\mathbb{N}$, $f\left(x_{1},\dots,x_{n}\right):\mathbb{N}^{n}\rightarrow\mathbb{N}$, we write $p\models f\left(x_{1},\dots,x_{n}\right)=0$ if for all $A\in\mathcal{U}$ there exists $a_{1},\dots,a_{n}\in A$ such that $f\left(a_{1},\dots,a_{n}\right)=0$.
Lemma 2.10.
[1, Lemma 2.1] Let $p\in\beta\mathbb{N}$, and let $f\left(x_{1},\dots,x_{n}\right),g\left(y_{1},\dots,y_{m}\right)$ be two functions such that $p\models\left(f_{1}\left(x_{1},\dots,x_{n}\right)=0\right)\wedge\left(g\left(y_{1},\dots,y_{m}\right)=0\right)$. Then $p\models\left(f_{1}\left(x_{1},\dots,x_{n}\right)=0\right)\wedge\left(g\left(y_{1},\dots,y_{m}\right)=0\right)\wedge\left(x_{1}=y_{1}\right)$.
As said, the above result allows us to recover the partition regularity of many equations and configurations. In fact, pick $p\in\overline{K(\beta\mathbb{N},\odot)}$ multiplicatively idempotent. Then:
[…]
  • as $p$ is idempotent, $p\models y_{1}\dot{}\dots\dot{y}_{m}=z$;
[…]
and so on, as we can always substitute a variable $x$ with a polynomial $P\left(y_{1},\dots,y_{n}\right)$ whenever $p\models x=P\left(y_{1},\dots,y_{n}\right)$. We refer to [1, Theorem 2.11] for a long list of such admissible substitutions.
Corollary 2.9 can be generalized also in another direction. Let us first fix a definition.
Definition 2.11.
We let
$EVdW=\{p\in\beta\mathbb{N}\mid\forall A\in p,\forall k\in\mathbb{N}\,\exists\,x,y\in A\ \text{s.t.}\ x^{y},\dots,x^{\left(y^{k}\right)}\in A\}.$
Theorem 2.12.
For all $\mathcal{U}\in\overline{K(\beta\mathbb{N},\odot)}$, for all $n\in\mathbb{N}_{>1}$, $n^{\mathcal{U}}\in EVdW$.
Actually, Theorem 2.12 can be seen as a very particular case of a much more general phenomenon involving $E_{1},E_{2}$.
Theorem 2.13.
Let $p,q\in\beta\mathbb{N}\setminus\{0\}$. Then:
  1. 1.
    If $q\in\overline{K(\beta\mathbb{N},\odot)}$ then $E_{1}\left(p,q\right)\in EVdW$. In particular, $E_{1}\left(p,q\right)$ witnesses the partition regularity of exponential triples;
  2. 2.
    If $q\in EVdW$ then $E_{2}(p,q)\in EVdW$. In particular, $E_{2}(p,q)$ witnesses the partition regularity of exponential triples.
Proof.
(1) Let $k\in\mathbb{N}$ and let $A\in E_{1}(p,q)$. We will show that $A$ contains a configuration of the form $x,y,x^{y},\dots,x^{\left(y^{k}\right)}$.
By definition, $I=\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid n^{m}\in A\}\in q\}\in p$. Let $n\in I$ and $J_{n}=\{m\in\mathbb{N}\mid n^{m}\in A\}$. Take $a,b\in J_{n}$ with $a,b,an^{b},\dots,an^{kb}\in J_{n}$, whose existence is ensured by Corollary 2.9 since, as $J_{n}\in q\in\overline{K(\beta\mathbb{N},\odot)}$, $J_{n}$ is multiplicatively piecewise syndetic set. Therefore,
$n^{a},n^{b},n^{an^{b}}=\left(n^{a}\right)^{\left(n^{b}\right)},\dots,n^{an^{kb}}=\left(n^{a}\right)^{\left(\left(n^{b}\right)^{k}\right)}\in A,$
and we conclude by letting $x=n^{a},y=n^{b}$.
(2) Let $q\in EVdW$, let $k\in\mathbb{N}$ be given and let $A\in E_{2}(p,q)$. By definition, $I=\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid m^{n}\in A\}\in q\}\in p$. Let $n\in I$ and let $J_{n}=\{m\in\mathbb{N}\mid m^{n}\in A\}$. Take $a,b\in J_{n}$ with $a,b,a^{\left(b^{n}\right)},\dots,a^{\left(b^{kn}\right)}\in J_{n}$. Hence
$a^{n},b^{n},\left(a^{\left(b^{n}\right)}\right)^{n}=a^{nb^{n}}=\left(a^{n}\right)^{\left(b^{n}\right)},\dots,\left(a^{\left(b^{kn}\right)}\right)^{n}=a^{nb^{kn}}=\left(a^{n}\right)^{\left(\left(b^{n}\right)^{k}\right)}\in A,$
and we conclude by letting $x=a^{n},y=b^{n}$.∎
As a consequence of [6, Theorem 6], if $Sc_{+}=\{p\in\beta\mathbb{N}\mid\forall A\in p\ \exists a,b\in A\ a+b\in A\}$, then $Sc_{+}\cap EVdW=\emptyset$; in particular,
$\overline{K(\beta\mathbb{N},\odot)}\cap EVdW=\emptyset,\overline{E(\oplus)}\cap EVdW=\emptyset,$
as $\overline{K(\beta\mathbb{N},\odot)}\cup\overline{E(\oplus)}\subseteq Sc_{+}$. Then, in particular, by Theorem 2.12 we also deduce that for all $n>1$, for all $p\in\overline{K(\beta\mathbb{N},\odot)}$, $n^{p}\notin\overline{K(\beta\mathbb{N},\odot)}$.
Non existence of exponential idempotents
As it is nowadays well-known, idempotents are often key ingredients in the search for infinite partition regular structures, the most famous example being the relationship between idempotent ultrafilters with respect to the sum and Hindman Theorem (see [5, Chapter 5]). As Examples 2.6, 2.7 suggest, having introduced the binary operations $E_{1},E_{2}$, it makes sense to study if they admit idempotents:
Definition 2.14.
We say that $p\in\beta\mathbb{N}$ is $E_{1}$-idempotent (resp. $E_{2}$-idempotent) if $E_{1}(p,p)=p$ (resp. if $E_{2}(p,p)=p$).
First of all, since $E_{1},E_{2}$ are not associative, we cannot approach the existence of idempotents problem via the usual straightforward application of Ellis’ Theorem (see [5, Theorem 2.5.] for details). Moreover, since $1$ is the only principal $E_{1}-$idempotent, and the only $E_{2}$-idempotent as well, from now on we will use “idempotent” to actually mean “nonprincipal idempotent”.
In the following we want to study properties of generalizations of the operation $(p,q)\rightarrow 2^{p}\odot q$, first introduced by Di Nasso and Ragosta in [2]. We will use the following notations: for $n\in\mathbb{N},p,q\in\beta\mathbb{N}$ we let:
$p\oleft_{n}q:=n^{p}\odot q;p\oright_{n}q:=p\odot n^{q}.$
The following result shows that the existence of idempotents for $E_{2}$ is related to that of ⊣⃝$\oleft_{n}⊣$-idempotents, whilst the existence of idempotents for $E_{1}$ is related to that of ⊢⃝$\oright_{n}⊢$-idempotents.
Proposition 2.15.
If $n\in\mathbb{N}_{>1}$, then $(1)\Rightarrow(2)\Rightarrow(3)$, where
  1. (i)
    there exists $p\in\beta\mathbb{N}$ such that $p\oplus n^{p}=p$;
  2. (ii)
    there exists $p\in\beta\mathbb{N}$ such that $p\oright_{n}p=p$;
  3. (iii)
    there exists $p\in\beta\mathbb{N}$ such that $E_{1}(p,p)=p$.
Analogously, $(4)\Rightarrow(5)\Rightarrow(6)$, where
  1. (iv)
    there exists $p\in\beta\mathbb{N}$ such that $n^{p}\oplus p=p$;
  2. (v)
    there exists $p\in\beta\mathbb{N}$ such that $p\oleft_{n}p=p$;
  3. (vi)
    there exists $p\in\beta\mathbb{N}$ such that $E_{2}(p,p)=p$.
Proof.
All the implications follows by taking $q$ as given by the hypothesis, by letting $p=n^{q}$ and by applying Proposition 2.4. For this reason, we prove only that $(v)$ entails $(vi)$, as the other implications are very similar.
$(v)\Rightarrow(vi)$ Let $q\in\beta\mathbb{N}$ be such that $q\oleft_{n}q=q$ and let $p=n^{q}$. Then $E_{2}(p,p)=E_{2}\left(n^{q},n^{q}\right)$. As $n^{q}=E_{2}(q,n)$, by Proposition 2.4.(6) we have that
$E_{2}\left(n^{q},n^{q}\right)=E_{2}\left(n^{q},E_{2}\left(q,n\right)\right)=E_{2}\left(n^{q}\odot q,n\right)=E_{2}\left(q,n\right)=p.\qed$
Moreover, in the search of $E_{1}$- and $E_{2}$-idempotent ultrafilters, we make the following observation:
Proposition 2.16.
Assume that there exist $p,q\in\beta\mathbb{N}\setminus\{0,1\}$ such that $p=p\odot E_{1}(q,p)$. Then $E_{1}(q,p)$ is $E_{1}$-idempotent.
Analogously, if there exist $p,q\in\beta\mathbb{N}\setminus\{0,1\}$ such that $p=E_{2}(p,q)\odot p$ then $E_{2}(p,q)$ is $E_{2}$-idempotent.
Proof.
We just have to apply Proposition 2.4.(5)-(6).
In fact, if $p=p\odot E_{1}(q,p)$ then
$E_{1}(q,p)=E_{1}(q,p)=E_{1}(q,p\odot E_{1}(q,p))=E_{1}\left(E_{1}(q,p),E_{1}(q,p)\right)$
and, if $p=E_{2}(p,q)\odot p$ then
$E_{2}(p,q)=E_{2}\left(E_{2}(p,q)\odot p,q\right)=E_{2}\left(E_{2}(p,q),E_{2}(p,q)\right).\qed$
All that has been said so far might point to the actual existence of exponential idempotents. Actually, as we are going to show, this is not the case for $E_{1}$, and at least not trivially the case for $E_{2}$.
As for $E_{1}$, the non existence of idempotents follows from an even stronger fact, that we now prove.
Theorem 2.17.
For all $p\in\beta\mathbb{N},q\in\beta\mathbb{N}\setminus\mathbb{N}$ we have that [^5] Notice, by contrast, that in Proposition 2.4 we proved that the apparently similar equation $E_{1}(p,q)=p$ has infinitely many solutions. $E_{1}(p,q)\neq q$; in particular, there are no $E_{1}-$idempotents ultrafilters in $\beta\mathbb{N}\setminus\mathbb{N}$.
Proof.
By [6, Lemma 23], there is a $4$-coloring $\mathbb{N}=A_{1}\cup\dots\cup A_{4}$ of $\mathbb{N}$ such that there is no monochromatic pair $b,a^{b}$ with $\log_{2}a\leq b$ and $a,b\geq 1$.
Now assume that $E_{1}(p,q)=q$; by definition, this means that for all $A\subseteq\mathbb{N}$ we have that
$A\in q\Leftrightarrow\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid n^{m}\in A\}\in q\}\in p.$
Now take $A_{i}\in\{A_{1},\dots,A_{4}\}$ such that $A_{i}\in q$. In particular, $B=\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid n^{m}\in A_{i}\}\in q\}\neq\emptyset$, so we can pick $n\in B$. Let $I_{n}:=\{m\in\mathbb{N}\mid n^{m}\in A_{i}\}$. As $q$ is non principal, $I_{n}$ must contain some $m>\log_{2}(n)$. But $m,n^{m}\in A_{i}$ with $\log_{2}(n)\leq m$, which is absurd. ∎
Putting together Theorem 2.17 and Propositions 2.15, 2.16, we get the following:
Corollary 2.18.
For all $n>1\in\mathbb{N}$ and $p,q\in\beta\mathbb{N}\setminus\{0,1\}$ we have that $p\oplus n^{p}\neq p$ and $p\odot E_{1}(q,p)\neq p$.
Notice that, in particular, also $p\oright_{n}{p}=p$ has no solution in $\beta\mathbb{N}$ if $n>1$.
Naively, from Theorem 2.17 it would be natural to imagine that some similar argument could be used to prove the following conjecture:
Conjecture 2.19.
$E_{2}$-idempotent ultrafilters do not exist.
Unfortunately, we have not been able to prove the above conjecture; we can, however, show a partial result towards its solution. The basic observation is the following: by Proposition 2.4, if $p$ is $E_{2}$-idempotent, we have that
$p=E_{2}(p,p)=E_{2}(p,E_{2}(p,p))=E_{2}(p\odot p,p).$
As, for all $n,m\in\mathbb{N},p\in\beta\mathbb{N}\setminus\mathbb{N}$ one has that $E_{2}(n,p)=E_{2}(m,p)\Leftrightarrow n=m$, it becomes natural to ask if it is possible to have $p\in\beta\mathbb{N}\setminus\mathbb{N}$ such that $p=p\odot p=E_{2}(p,p)$; the answer is no.
Theorem 2.20.
There is no ultrafilter $p\in\beta\mathbb{N}\setminus\mathbb{N}$ that is multiplicatively and $E_{2}$-idempotent.
Proof.
By contrast, let $p$ be such an ultrafilter. Let $S:\mathbb{N}\rightarrow\mathbb{N}$ be the function that maps $n=p_{1}^{a_{1}}\cdot\dots\cdot p_{k}^{a_{k}}$ to $\sum_{i\leq k}a_{i}$, where $n=p_{1}^{a_{1}}\cdot\dots\cdot p_{k}^{a_{k}}$ is the prime decomposition of $n$. Observe that $S(E_{2}(p_{1},p_{2}))=p_{1}\odot S(p_{2})$ and $S(p_{1}\odot p_{2})=S(p_{1})\oplus S(p_{2})$ for all $p_{1},p_{2}\in\beta\mathbb{N}$. In particular, $S(p)=S(E_{2}(p,p))=p\odot S(p)$ and $S(p)=S(p\odot p)=S(p)\oplus S(p)$, therefore $S(p)$ is additively idempotent and $p\odot S(p)=S(p)\oplus S(p)$. If $S(p)\in\mathbb{N}$, then by additive idempotency we would get that $S(p)=0$, from which $p=1$, which is absurd as $p\in\beta\mathbb{N}\setminus\mathbb{N}$. So $S(p)$ is additively idempotent and nonprincipal, so $n\mathbb{N}\in S(p)$ for all $n\in\mathbb{N}$. Therefore we can apply [5, Theorem 13.14] to deduce that it must be $p\odot S(p)\neq S(p)\oplus S(p)$, reaching the desired contrast. ∎
3   Open questions
In this section, we summarize some questions which we could not solve, as well as some possible directions we believe it could be interesting to pursue in the future.
The first question originates from the fact that, in [6], Sahasrabudhe proved the partition regularity of the configuration $\{x,y,xy,x^{y}\}$. We tried to recover a more direct proof of this fact, along the lines of what we developed in Section 2.1, but we failed. Our guess is that, by choosing $p$ in a reasonable way, the partition regularity of $\{x,y,xy,x^{y}\}$ should be witnessed by $2^{p}$. We conjecture that a family of ultrafilters $p$ that should make our claim true is the family of those $p\in E\left(K(\beta\mathbb{N},\odot)\right)$ such that each member of $p$ is also additive central. Summarizing, our first question is the following:
Question 3.1.
Let $p\in E\left(K(\beta\mathbb{N},\odot)\right)$ be such that each member of $p$ is also additive central. Is it true or not that each $A\in p$ contains a configuration of the form $\{m,n,m+n,m2^{n}\}$?
Notice that an affirmative answer to the above question would immediately entail, by using the same methods adopted in this paper, that each member of $2^{p}$ contains patterns of the form $\{x,y,xy,x^{y}\}$.
The following question regards several technical properties of the exponential operations $E_{1},E_{2}$. We proved the existence and non-existence of ultrafilters satisfying certain equalities; for some other similar ones, we did not manage to find a definitive answer.
Question 3.2.
Are there $p,q\in\beta\mathbb{N}\setminus\mathbb{N}$ that solve any of the following equations
  1. 1.
    $E_{1}(p,q)=E_{1}(q,p)$;
  2. 2.
    $E_{2}(p,q)=E_{2}(q,p)$;
  3. 3.
    $E_{1}(p,q)=E_{2}(q,p)$;
  4. 4.
    $E_{1}(q,p)\odot p=p$;
  5. 5.
    $E_{1}(q,p)\odot q=q$;
  6. 6.
    $E_{2}(q,p)\odot p=p$;
  7. 7.
    $E_{2}(q,p)\odot q=q$.
Of course, as discussed in the paper, more relevant questions are the following:
Question 3.3.
Does it exist $p$ that is $E_{2}$-idempotent? Does it exist $p$ that is ⊣⃝$\oleft_{n}⊣$-idempotent?
Finally, as a general direction connected to all the above questions, we believe that it would be interesting to further study the properties of exponentially rich ultrafilters; of course, this rather vague idea could be made precise in several different ways. We believe it would be interesting to start by studying the following:
Question 3.4.
What are the algebraical properties of $EVdW$, besides Theorem 2.12?

### v4

@@
 1   Introduction
-A large class of problems in arithemtic Ramsey theory regards the so-called partition regular configurations and equations on =$\mathbb{N}=\{1,2,\dots\}$.
+A large class of problems in arithmetic Ramsey theory regards the so-called partition regular configurations and equations on =$\mathbb{N}=\{1,2,\ldots\}$.
 Definition 1.1.
-Let $f_{1}\left(x_{1},\dots,x_{n}\right),\dots,f_{m}\left(x_{1},\dots,x_{n}\right):\mathbb{N}^{n}\rightarrow\mathbb{N}$.
-We say that the configuration $\{f_{1}\left(x_{1},\dots,x_{n}\right),\dots,f_{m}\left(x_{1},\dots,x_{n}\right)\}$ is partition regular on $\mathbb{N}$ (PR from now on) if for all finite partitions $\mathbb{N}=A_{1}\cup\dots\cup A_{k}$ there are $i\leq k$ and $x_{1},\dots,x_{n}\in\mathbb{N}$ such that for all $j\leq m$ $f_{j}\left(x_{1},\dots,x_{n}\right)\in A_{i}$.
-Analogously, we say that the system $\sigma\left(x_{1},\dots,x_{n}\right)$ defined by
-$\begin{cases}f_{1}\left(x_{1},\dots,x_{n}\right)=0,\\ \ \ \ \ \ \ \ \ \ \vdots\\ f_{m}\left(x_{1},\dots,x_{n}\right)=0\end{cases}$
-is PR if for all finite partitions $\mathbb{N}=A_{1}\cup\dots\cup A_{k}$ there are $i\leq k$ and $x_{1},\dots,x_{n}\in A_{i}$ such that $\sigma\left(x_{1},\dots,x_{n}\right)=0$.
+Let $f_{1}\left(x_{1},\ldots,x_{n}\right),\ldots,f_{m}\left(x_{1},\ldots,x_{n}\right):\mathbb{N}^{n}\rightarrow\mathbb{N}$.
+We say that the configuration $\{f_{1}\left(x_{1},\ldots,x_{n}\right),\ldots,f_{m}\left(x_{1},\ldots,x_{n}\right)\}$ is partition regular on $\mathbb{N}$ (PR from now on) if for all finite partitions $\mathbb{N}=A_{1}\cup\ldots\cup A_{k}$ there are $i\leq k$ and $x_{1},\ldots,x_{n}\in\mathbb{N}$ such that for all $j\leq m$ $f_{j}\left(x_{1},\ldots,x_{n}\right)\in A_{i}$.
+Analogously, we say that the system $\sigma\left(x_{1},\ldots,x_{n}\right)$ defined by
+$\begin{cases}f_{1}\left(x_{1},\ldots,x_{n}\right)=0,\\ \ \ \ \ \ \ \ \ \ \vdots\\ f_{m}\left(x_{1},\ldots,x_{n}\right)=0\end{cases}$
+is PR if for all finite partitions $\mathbb{N}=A_{1}\cup\ldots\cup A_{k}$ there are $i\leq k$ and $x_{1},\ldots,x_{n}\in A_{i}$ such that $\sigma\left(x_{1},\ldots,x_{n}\right)=0$.
 In [8], Schur proved one of the first and most important results in Arithmetic Ramsey Theory.
-Theorem 1.2 (Additive Schur Theorem).
+Theorem 1.2 (Additive Schur’s Theorem).
 The configuration $\{x,y,x+y\}$ is PR.
-Equivalently, the additive Schur Theorem states that the equation $x+y=z$ is PR.
+Equivalently, the additive Schur’s Theorem states that the equation $x+y=z$ is PR.
 Applying transformations to partitions allows to easily deduce new results by just considering the shape of transformed partition regular configurations.
-A useful trick, that sometimes allows turning additive structures in multiplicative structures, and multiplicative structures in exponential ones, is that of considering $\log_{2}$-base partitions: for $B\subseteq\mathbb{N}$, let $2^{-B}=\{n\in\mathbb{N}\mid 2^{n}\in B\}$ and, given a partition $\mathbb{N}=A_{1}\cup\dots\cup A_{k}$, let the associated $\log_{2}$-base partition be given by $\mathbb{N}=2^{-A_{1}}\cup\dots\cup 2^{-A_{k}}$.
-For example, given any finite partition of $\mathbb{N}$, by Theorem 1.2 in the $\log_{2}$-base partition we find monochromatic $x,y,z$ with $x+y=z$; hence, $2^{x},2^{y},2^{z}=2^{x+y}=2^{x}\cdot 2^{y}$ were monochromatic with respect to the original partition.
+A useful trick, that sometimes allows turning additive structures in multiplicative structures, and multiplicative structures in exponential ones, is that of considering $\log_{2}$-base partitions: for $B\subseteq\mathbb{N}$, let $2^{-B}=\{n\in\mathbb{N}\mid 2^{n}\in B\}$ and, given a partition $\mathbb{N}=A_{1}\cup\ldots\cup A_{k}$, let the associated $\log_{2}$-base partition be given by $\mathbb{N}=2^{-A_{1}}\cup\ldots\cup 2^{-A_{k}}$.
+For example, given any finite partition of $\mathbb{N}$, in the $\log_{2}$-base partition we find monochromatic $x,y,z$ with $x+y=z$ by Theorem 1.2; hence, $2^{x},2^{y},2^{z}=2^{x+y}=2^{x}\cdot 2^{y}$ were monochromatic with respect to the original partition.
 This proves the multiplicative version of Schur.
-Theorem 1.3 (Multiplicative Schur Theorem).
+Theorem 1.3 (Multiplicative Schur’s Theorem).
 The configuration $\{x,y,x\cdot y\}$ is PR.
 Equivalently, the multiplicative Schur Theorem states that the equation $xy=z$ is PR.
-Recently, in [8], the question about the partition regularity of $\{x,y,x^{y}\}$, which we will call an exponential triple, was posed and partially answered, by using ultrafilters, by Sisto; its full solution was first given in [6] by Sahasrabudhe who, among other generalizations, proved the exponential Schur Theorem [^2] At the best of my knowledge, this theorem has not been called explicitly like this in the literature before.:
-Theorem 1.4 (Exponential Schur Theorem).
+Recently, in [9], the question about the partition regularity of $\{x,y,x^{y}\}$, which we will call an exponential triple, was posed and partially answered, by using ultrafilters, by Sisto; its full solution was first given in [7] by Sahasrabudhe who, among other generalizations, proved the exponential Schur’s Theorem [^2] At the best of our knowledge, this theorem has not been called explicitly like this in the literature before.:
+Theorem 1.4 (Exponential Schur’s Theorem).
 The configuration $\{x,y,x^{y}\}$ is PR.
 Again, the above theorem can be rephrased in terms of equations: it states that $x^{y}=z$ is PR.
-Sahasrabudhe’s proof used repeated applications of van der Waerden theorem and the compactness property on the convergence of subsequences in the space of $k$-partitions of $\mathbb{N}$.
-One of the key ingredients in his proof was the use of the function $f(n,m):=2^{n}m$, combined with associated $\log_{2}$-base partitions.
-Starting from observations about this function, in [2, Theorem 1.1] Di Nasso and Ragosta found an ultrafilters proof of the partition regularity of the configuration $\{x,y,x2^{y}\}$, from which the exponential Schur’s theorem can be immediately deduced by the same $\log_{2}$-partition trick that allows proving the multiplicative version of Schur Theorem from the additive one (see [2, Corollary 1.3]).
-Their proof was based on a clever trick based on tensor products of ultrafilters [^3] Through the paper we assume the reader to have a basic knowledge of the main properties of $\beta\mathbb{N}$; we refer to the monograph [5] for an expensive presentation of ultrafilters and their applications to combinatorics.
-and the extension to $\left(\beta\mathbb{N}\right)^{2}$ of the same function $f(n,m)=2^{n}m$ used by Sahasrabudhe is applied.
+Sahasrabudhe’s proof used the compactness property on the convergence of subsequences in the space of $k$-partitions of $\mathbb{N}$ and repeated applications of van der Waerden’s Theorem:
+Theorem 1.5 (Van der Waerden’s Theorem).
+For all $k\in\mathbb{N}$ the configuration $\{x,x+y,\ldots,x+ky\}$ is PR.
+One of the key ingredients in Sahasrabudhe’s proof was the use of the function $f(n,m):=2^{n}m$, combined with associated $\log_{2}$-base partitions.
+Starting from observations about this function, in [2, Theorem 2.1] Di Nasso and Ragosta gave a very simple and elegant ultrafilters proof of the partition regularity of the configuration $\{x,y,2^{x}y\}$, from which the exponential Schur’s theorem can be immediately deduced by the same $\log_{2}$-partition trick that allows proving the multiplicative Schur’s Theorem from the additive one (see [2, Corollary 2.3]).
+Their proof was based on tensor products of ultrafilters [^3] We will recall some basic facts and definitions regarding ultrafilters in Section 2.1, but we assume the reader to have a basic knowledge of the main properties of $\beta\mathbb{N}$; we refer to the monograph [5] for a detailed presentation of ultrafilters and their applications to combinatorics.
+and the extension to $\left(\beta\mathbb{N}\right)^{2}$ of the same function $f(n,m)=2^{n}m$ used by Sahasrabudhe.
 Di Nasso and Ragosta used ultrafilters methods also to study the partition regularity of infinite exponential configurations in [3].
-As its often the case with partition regularity problems, the ultrafilters approach provides a simple and informative proof that allows for several generalizations: this is a common situation, as proven by an enormous literature on the subject (instead of citing here hundreds of papers, we refer to [5] and its bibliography).
+As it is often the case with partition regularity problems, the ultrafilters proof allows for several generalizations: this is a common situation, as proven by an enormous literature on the subject (instead of citing here hundreds of papers, we refer to [5] and its bibliography).
 Our main goal is to start a systematic study of which ultrafilters are related to exponential triples, and what other combinatorial properties such ultrafilters have.
-This will be done by introducing two different exponentiations of ultrafilters, and by studying their properties.
-In particular, we will prove that all multiplicatively piecewise syndetic sets contain exponential triples which, combined with our exponentiations of ultrafilters, will allow to show that large families of ultrafilters witness the exponential Schur Theorem, in the sense that all their sets contain exponential triples.
-When a binary operations on $\beta\mathbb{N}$ is introduced, it is natural to search for idempotents since, as it is well-known since Galvin and Glazer’s proof of Hindman Theorem, such idempotents have plenty of interesting combinatorial properties.
-Henceforth, we study and solve negatively (completely for one extension, partially for the other) the problem of the existence of idempotents in Section 2.
-It turns out that idempotents for such exponentiations are strictly related to idempotents for other exponential-like operations, namely generalizations of the operation $(p,q)\rightarrow 2^{p}\odot q$, which first appeared in [2].
-The novelty of most of our results generates a large number of open questions and future possible directions. We conclude the paper with a discussion of many of these in Section 3.
+This will be done by introducing, in Section 2, two different exponentiations of ultrafilters, named $E_{1},E_{2}$, and by studying their algebraic properties and their connections with the exponential Schur’s Theorem.
+In Section 2.3 we study the existence of idempotents for $E_{1},E_{2}$, and prove that $E_{1}$ does not admit nonprincipal idempotents, whilst $E_{2}$ does not admit nonprincipal idempotents that are simultaneously multiplicative idempotents.
+The novelty of most of our results generates a large number of open questions and future possible directions.
+We conclude the paper with a discussion of many of these in Section 3.
 2   Exponentiations of ultrafilters
+Preliminaries
 As ultrafilters are the key notion of this paper, we assume the reader to have a basic knowledge of the algebra of $\beta\mathbb{N}$.
-We just briefly recall a few definitions, mostly so that we can fix a few notations.
-A notion that we will use often is that of tensor product of ultrafilters in $\beta\mathbb{N}$, defined as follows:
+We just briefly recall a few definitions, mostly so that we can fix some notations.
+Sums and products of ultrafilters will be denoted by $\oplus$,$\odot$ , ⊙ respectively.
+Those notion are related to that of tensor product of ultrafilters in $\beta\mathbb{N}$, which is defined as follows:
 Definition 2.1.
@@
 Moreover, we will need to talk about extensions of functions to $\beta\mathbb{N}$: for $f:\mathbb{N}^{n}\rightarrow\mathbb{N}$, we will denote by $\widehat{f}$ its unique continuous extension $\widehat{f}:\beta(\mathbb{N}^{n})\rightarrow\beta\mathbb{N}$.
+For example, when $S(x,y)=x+y$ is the sum, we have that $\hat{S}(p\odot q)=p\oplus q$ for all $p,q\in\beta\mathbb{N}$.
 As usual, to simplify notations we identify each natural number $n\in\mathbb{N}$ with the principal idempotent it generates, so that $\mathbb{N}\subseteq\beta\mathbb{N}$.
-The only purely combinatorial notion that we will need is that of multiplicatively piecewise set, defined as follows.
-A set $A\subseteq\mathbb{N}$ is multiplicatively syndetic if there exists a finite set $F$ such that for all $n\in\mathbb{N}$ there exists $f\in F$ such that $fx\in A$; $A$ is multiplicatively thick if for all $F$ finite subset of $\mathbb{N}$ there exists $x\in\mathbb{N}$ such that $xF\subseteq A$; $A$ is multiplicatively piecewise syndetic if $A=B\cap C$ where $B$ is multiplicatively syndetic and $C$ is multiplicatively thick.
-The relevant property of multiplicatively piecewise sets for us is the following:
-Proposition 2.2.
-[5, Theorem 4.40] $A\subseteq\mathbb{N}$ is multiplicatively piecewise syndetic if and only if there exists $p\in K(\beta\mathbb{N},\odot)$ such that $A\in p$.
+As well-known, ultrafilters are often a fundamental tool to study PR problems. For configurations, the connection between ultrafilters and partition regularity reads as follows:
+Fact 2.2.
+Let $f_{1}\left(x_{1},\ldots,x_{n}\right),\ldots,f_{m}\left(x_{1},\ldots,x_{n}\right):\mathbb{N}^{n}\rightarrow\mathbb{N}$. The following are equivalent:
+  1. 1.
+    $\{f_{1}\left(x_{1},\ldots,x_{n}\right),\ldots,f_{m}\left(x_{1},\ldots,x_{n}\right)\}$ is PR;
+  2. 2.
+    $\exists p\in\beta\mathbb{N}\ \forall A\in p\ \exists x_{1},\ldots,x_{n}\in\mathbb{N}\ f_{1}\left(x_{1},\ldots,x_{n}\right),\ldots,f_{m}\left(x_{1},\ldots,x_{n}\right)\in A$.
+The above fact leads to the following definition:
+Definition 2.3.
+Let $p\in\beta\mathbb{N}$ and let $f_{1}\left(x_{1},\ldots,x_{n}\right),\ldots,f_{m}\left(x_{1},\ldots,x_{n}\right):\mathbb{N}^{n}\rightarrow\mathbb{N}$.
+We say that $p$ witnesses the partition regularity of the configuration $\{f_{1}\left(x_{1},\ldots,x_{n}\right),\ldots,f_{m}\left(x_{1},\ldots,x_{n}\right)\}$ if for all $A\in p$ there exist $x_{1},\ldots,x_{n}\in A$ such that $f\left(x_{1},\ldots,x_{n}\right)=0$.
+In this case, we write $p\models\{f_{1}\left(x_{1},\ldots,x_{n}\right),\ldots,f_{m}\left(x_{1},\ldots,x_{n}\right)\}$.
+In the next sections we will use a few specific results from other papers.
+For the readers easiness, we recall them here.
+The first states that when the PR of different systems of equations are witnessed by the same ultrafilter, these configurations can be mixed.
+Lemma 2.4.
+[1, Lemma 2.1] Let $p\in\beta\mathbb{N}$, and let $f\left(x_{1},\ldots,x_{n}\right),g\left(y_{1},\ldots,y_{m}\right)$ be two functions such that $p\models\left(f_{1}\left(x_{1},\ldots,x_{n}\right)=0\right)\wedge\left(g\left(y_{1},\ldots,y_{m}\right)=0\right)$.
+Then $p\models\left(f_{1}\left(x_{1},\ldots,x_{n}\right)=0\right)\wedge\left(g\left(y_{1},\ldots,y_{m}\right)=0\right)\wedge\left(x_{1}=y_{1}\right)$.
+The second result that we will need is [4, Theorem 1.6], which states that multiplicatively piecewise syndetic sets contains some exponential structure.
+Theorem 2.5.
+[4, Theorem 1.6] Let $n\in\mathbb{N},n>1$ and let $A\subseteq\mathbb{N}$ be multiplicatively piecewise syndetic. Then there exists $x,y\in A$ such that $xn^{y}\in A$.
+The third result we need is [6, Theorem A], that states that certain ultrafilter equations admits no solution.
+Theorem 2.6.
+[6, Theorem A] Let $p\in\beta\mathbb{N}\setminus\mathbb{N}$, $a,b\in\mathbb{N}$ and $u,v\in\beta\mathbb{N}$. Suppose that $u\oplus ap=v\oplus bp$. Then $a=b$.
+Finally, we will the non-PR of combined sums and exponentiations stated in [7, Theorem 6].
+Theorem 2.7.
+[7, Theorem 6] The configuration $\{x,y,x^{y},a,b,a+b\}$ is not PR.
 Two extensions of exponentiation to ultrafilters
@@
 $A\in p\oplus q\Leftrightarrow\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid n+m\in A\}\in q\}\in p,$
-and the reverse sum $p\tilde{\oplus}q:=q\oplus p$; if $S(n,m):=n+m$ denotes the sum of natural numbers, then $p\oplus q=\widehat{S}(p\otimes q)$ and $q\oplus p=\widehat{S}(p\otimes q)$.
+and the reverse sum $p\tilde{\oplus}q:=q\oplus p$; if $S(n,m):=n+m$ denotes the sum of natural numbers, then $p\oplus q=\widehat{S}(p\otimes q)$ and $q\oplus p=\widehat{S}(q\otimes p)$.
 Whilst different, $\oplus$ and $\tilde{\oplus}$ have the same properties, modulo properly translating their meaning by switching “left” with “right” notions.
@@
 Given $p,q\in\beta\mathbb{N}$ we let
-$E_{1}(p,q):=\{A\subseteq\mathbb{N}\mid\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid n^{m}\in q\}\in p\},$
+$E_{1}(p,q):=\{A\subseteq\mathbb{N}\mid\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid n^{m}\in A\}\in q\}\in p\},$
 and
-$E_{2}(p,q):=\{A\subseteq\mathbb{N}\mid\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid m^{n}\in q\}\in p\}.$
-Equivalently, if $e:\mathbb{N}^{2}\rightarrow\mathbb{N}$ is the exponentiation $e(n,m)=n^{m}$, then $E_{1}(p,q)=\widehat{e}(p\otimes q)$, whilst $E_{2}(p,q)=\widehat{e}(q\otimes p)$.
-For $n,m\in\mathbb{N}$, trivially we have that $E_{1}(n,m)=E_{2}(m,n)=n^{m}$; more in general, for $n\in\mathbb{N},p\in\beta\mathbb{N}$ we have that $E_{1}(n,p)=E_{2}(p,n):=p^{n}$ and $E_{1}(p,n)=E_{2}(n,p)=n^{p}$ as, for all $A\subseteq\mathbb{N}$
+$E_{2}(p,q):=\{A\subseteq\mathbb{N}\mid\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid m^{n}\in A\}\in q\}\in p\}.$
+Equivalently, if $e_{1},e_{2}:\mathbb{N}^{2}\rightarrow\mathbb{N}$ are the exponentiations $e_{1}(n,m)=n^{m},e_{2}(n,m)=m^{n}$, then $E_{1}(p,q)=\widehat{e_{1}}(p\otimes q)$ and $E_{2}(p,q)=\widehat{e_{2}}(p\otimes q)$.
+For $n,m\in\mathbb{N}$, trivially we have that $E_{1}(n,m)=E_{2}(m,n)=n^{m}$; more in general, for $n\in\mathbb{N},p\in\beta\mathbb{N}$ we have that $E_{1}(n,p)=E_{2}(p,n)$ and $E_{1}(p,n)=E_{2}(n,p)$ as, for all $A\subseteq\mathbb{N}$
 $A\in E_{1}(n,p)\Leftrightarrow\{m\in\mathbb{N}\mid n^{m}\in A\}\in p\Leftrightarrow A\in E_{2}(p,n),$
 $A\in E_{1}(p,n)\Leftrightarrow\{m\in\mathbb{N}\mid m^{n}\in A\}\in p\Leftrightarrow A\in E_{2}(n,p).$
-Again, in terms of extension of functions, the above shows that, for $n$ fixed, $p^{n}=\widehat{e_{n}}(p)$, where $e_{n}:\mathbb{N}\rightarrow\mathbb{N}$ is the function that maps $m$ to $m^{n}$, and $n^{p}=\widehat{e^{\prime}_{n}}(p)$, where $e^{\prime}_{n}:\mathbb{N}\rightarrow\mathbb{N}$ is the function that maps $m$ to $n^{m}$.
-Notation: we will continue to use the simplified notation $n^{p}$ to denote any of $E_{1}(n,p),E_{2}(p,n)$, and we will use the simplified notation $p^{n}$ to denote any of $E_{1}(p,n),E_{2}(n,p)$.
+Notation: we will use the simplified notation $n^{p}$ to denote any of $E_{1}(n,p),E_{2}(p,n)$, and we will use the simplified notation $p^{n}$ to denote any of $E_{1}(p,n),E_{2}(n,p)$.
 At first sight, it might look like the equality $E_{1}(p,q)=E_{2}(q,p)$ could be true for all $p,q\in\beta\mathbb{N}$; however, this is not the case.
-In fact, for example, let both $p,q$ contain the set $P$ of primes.
-Then
+Example 2.9.
+Let both $p,q$ be nonprincipal ultrafilters that contain the set $P$ of primes. Then
 $P_{1}=\{a^{b}\mid a,b\in P,a<b\}\in E_{1}(p,q),P_{2}=\{a^{b}\mid a,b\in P,a>b\}\in E_{2}(p,q)$
-but $P_{1}\cap P_{2}=\emptyset$, so $E_{1}(p,q)\neq E_{2}(q,p)$. We will provide some examples of nonprincipal ultrafilters $p,q$ satisfying $E_{1}(p,q)=E_{2}(q,p)$ after Proposition 2.4.
-$E_{1}$ and $E_{2}$ allow to extend, in a proper way, several algebraical properties of exponentiation from natural numbers to ultrafilters.
-We list here those that we will need in the rest of the paper.
-To avoid cumbersome reformulations, some of the properties proven in the following proposition are stated for rational numbers and ultrafilters on the rationals, to which the exponentiation functions $E_{1},E_{2}$ can be extended trivially.
+but $P_{1}\cap P_{2}=\emptyset$, so $E_{1}(p,q)\neq E_{2}(q,p)$.
+We will provide some examples of nonprincipal ultrafilters $p,q$ satisfying $E_{1}(p,q)=E_{2}(q,p)$ after Proposition 2.10.
+To prove this proposition, it is helpful to use fractional powers of ultrafilters; to make proper sense of these powers, we prefer to state and prove Proposition 2.10 for ultrafilters in $\beta\mathbb{R}^{+}_{d}$, namely elements of the Stone-Čech compactification of +:=$\mathbb{R}^{+}:=\{x\in\mathbb{R}\mid x>0\}$ endowed with the discrete topology.
+We extend the exponentiations to $\beta\mathbb{R}_{d}$ as follows: given $p,q\in\beta\mathbb{R}^{+}_{d}$, we let:
+$E_{1}(p,q):=\{A\subseteq\mathbb{R}^{+}\mid\{x\in\mathbb{R}^{+}\mid\{y\in\mathbb{R}^{+}\mid x^{y}\in A\}\in q\}\in p\},$
+and
+$E_{2}(p,q):=\{A\subseteq\mathbb{R}^{+}\mid\{x\in\mathbb{R}^{+}\mid\{y\in\mathbb{R}^{+}\mid y^{x}\in A\}\in q\}\in p\}.$
+For $r\in\mathbb{R}^{+}$ and $p\in\beta\mathbb{R}^{+}_{d}$, we will keep the simplified notation $r^{p}$ to denote any of $E_{1}(r,p),E_{2}(p,r)$ (which are equal), and $p^{r}$ to denote any of $E_{1}(p,r),E_{2}(r,p)$ (which are also equal).
 Proposition 2.10.
-Let $p,q,r\in\beta\mathbb{N}\setminus\{0,1\},n,m\in\mathbb{N}\setminus\{1\},a\in\mathbb{Q}\setminus\{0\}$. The following properties hold:
+Let $p,q,r\in\beta\mathbb{R}_{d}^{+}\setminus\{1\}$. Let $a,b\in\mathbb{R}^{+}$ with $a\geq b>1$. The following properties hold:
   1. 1.
@@
   2. 2.
-    $n^{p}=n^{q}\Leftrightarrow p=q$ and $p^{n}=q^{n}\Leftrightarrow p=q$;
+    $a^{p}=b^{q}\Leftrightarrow\left(\log_{a}(b)\right)p=q$;
   3. 3.
-    $n^{p}=m^{p}\Leftrightarrow n=m$ and $p^{n}=p^{m}\Leftrightarrow n=m$;
+    $p^{a}=q^{b}\Leftrightarrow p^{\frac{a}{b}}=q$;
   4. 4.
-    $n^{p}\odot n^{q}=n^{p\oplus q}$, but $E_{1}(r,p)\odot E_{1}(r,q)\neq E_{1}(r,p\oplus q)$ and $E_{2}(p,r)\odot E_{2}(q,r)\neq E_{2}(p\oplus q,r)$ in general;
+    $a^{p}\odot a^{q}=a^{p\oplus q}$;
   5. 5.
@@
   6. 6.
-    $E_{2}\left(p,E_{2}(q,r)\right)=E_{2}\left(p\odot q,r\right)$; in particular, if $p$ is multiplicatively idempotent, $E_{2}\left(p,E_{2}(p,q)\right)=E_{2}(p,q)$.
+    $E_{2}\left(p\odot q,r\right)=E_{2}\left(p,E_{2}(q,r)\right)$; in particular, if $p$ is multiplicatively idempotent, $E_{2}(p,q)=E_{2}\left(p,E_{2}(p,q)\right)$.
 Proof.
-All these proofs consist, basically, of straightforward checks on the definitions.
-$(1)$ Let $A\subseteq\mathbb{N}$. Then
-$A\in E_{1}\left(p^{a},q\right)\Leftrightarrow\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid n^{m}\in A\}\in q\}\in p^{a}\Leftrightarrow$
-$\{n\in\mathbb{N}\mid n^{a}\in\{l\in\mathbb{N}\mid\{m\in\mathbb{N}\mid l^{m}\in A\}\in q\}\}\in p\Leftrightarrow\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid\left(n^{a}\right)^{m}\in A\}\in q\}\in p\Leftrightarrow$
-$\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid\left(n^{am}\right)\in A\}\in q\}\in p\Leftrightarrow\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid\left(n^{m}\right)\in A\}\in aq\}\in p\Leftrightarrow$
-$A\in E_{1}\left(p,aq\right).$
+$(1)$ Let $A\subseteq\mathbb{R}^{+}$. Then
+$A\in E_{1}\left(p^{a},q\right)\Leftrightarrow\{x\in\mathbb{R}^{+}\mid\{y\in\mathbb{R}^{+}\mid x^{y}\in A\}\in q\}\in p^{a}\Leftrightarrow$
+$\{z\in\mathbb{R}^{+}\mid z^{a}\in\{x\in\mathbb{R}^{+}\mid\{y\in\mathbb{R}^{+}\mid x^{y}\in A\}\in q\}\}\in p\Leftrightarrow$
+$\{z\in\mathbb{R}^{+}\mid\{y\in\mathbb{R}^{+}\mid\left(z^{a}\right)^{y}\in A\}\in q\}\in p\Leftrightarrow\{z\in\mathbb{R}^{+}\mid\{y\in\mathbb{R}^{+}\mid z^{\left(ay\right)}\in A\}\in q\}\in p\Leftrightarrow$
+$\{z\in\mathbb{R}^{+}\mid\{t\in\mathbb{R}^{+}\mid z^{t}\in A\}\in aq\}\in p\Leftrightarrow A\in E_{1}(p,aq).$
 The $E_{2}$-case is analogous.
-$(2)$ If $p=q$, both equivalences are trivial.
-If $n^{p}=n^{q}$ it means that for all $A\subseteq\mathbb{N}$, if $B=\{a\in\mathbb{N}\mid n^{a}\in A\}$ then $B\in p\Leftrightarrow B\in q$, which shows that if we let $P_{n}=\{a\in\mathbb{N}\mid\exists b\in\mathbb{N}\ a=b^{n}\}$ and $log_{n}:P_{n}\rightarrow\mathbb{N}$ be the $n$-base logarithm and $\widehat{log_{n}}$ be its extension to ultrafilters containing $P_{n}$ then $p=\widehat{log_{n}}\left(n^{p}\right)=\widehat{log_{n}}\left(n^{q}\right)=q$.
-The other equivalence can be proven similarly.
-$(3)$ If $n=m$, both equivalences are trivial.
-If $n^{p}=m^{p}$, first we observe that $n^{\mathbb{N}}=\{n^{k}\mid k\in\mathbb{N}\}\in n^{p}$ and $m^{\mathbb{N}}=\{m^{k}\mid k\in\mathbb{N}\}\in m^{p}$.
-Therefore, if $n^{p}=m^{p}$ we have that $n^{\mathbb{N}}\cap m^{\mathbb{N}}\in n^{p}$.
-Hence $n^{\mathbb{N}}\cap m^{\mathbb{N}}\neq\emptyset$, which forces $n^{a}=m$ for some $a\in\mathbb{Q}$, so $n^{p}=n^{ap}$ from $(1)$, which from $(2)$ forces $p=ap$, which forces $a=1$, so $n=m$.
-The other equivalence is similar.
-$(4)$ First, let us show that $n^{p}\odot n^{q}=n^{p\oplus q}$: let $A\subseteq\mathbb{N}$. Then
-$A\in n^{p}\odot n^{q}\Leftrightarrow\{a\in\mathbb{N}\mid\{b\in\mathbb{N}\mid ab\in A\}\in n^{q}\}\in n^{p}\Leftrightarrow$
-$\{d\in\mathbb{N}\mid n^{d}\in\{a\in\mathbb{N}\mid\{c\in\mathbb{N}\mid n^{c}\in\{b\in\mathbb{N}\mid ab\in A\}\in q\}\}\}\in p\Leftrightarrow$
-$\{d\in\mathbb{N}\mid\{c\in\mathbb{N}\mid n^{c}\cdot n^{d}\in A\}\in q\}\in p\Leftrightarrow\{d\in\mathbb{N}\mid\{c\in\mathbb{N}\mid n^{c+d}\in A\}\in q\}\in p\Leftrightarrow$
-$\{e\in\mathbb{N}\mid n^{e}\in A\}\in p\oplus q\Leftrightarrow A\in n^{p\oplus q}.$
-For the other claims: let $P=\{n\in\mathbb{N}\mid n\ \text{is a prime}\}$ and let $P\in r$. Let $p,q$ be nonprincipal. Then a direct computation shows that
+$(2)$ $\Rightarrow$)⇒ ) Consider the function $\log_{a}:\mathbb{R}^{+}\rightarrow\mathbb{R}$.
+A straightforward computation shows that $\widehat{\log_{a}}\left(a^{p}\right)=p$, hence our hypothesis entails that $p=\widehat{\log_{a}}\left(b^{q}\right)$.
+Now we observe that, for any $A\subseteq\mathbb{R}^{+}$,
+$A\in\widehat{\log_{a}}\left(b^{q}\right)\Leftrightarrow(\log_{a})^{-1}(A)\in\left(b^{q}\right)\Leftrightarrow\{x\in\mathbb{R}^{+}\mid\{y\in\mathbb{R}^{+}\mid x^{y}\in(\log_{a})^{-1}(A)\}\in q\}\in b\Leftrightarrow$
+$\{y\in\mathbb{R}^{+}\mid b^{y}\in(\log_{a})^{-1}(A)\}\in q\Leftrightarrow\{y\in\mathbb{R}^{+}\mid\log_{a}\left(b^{y}\right)\in A\}\in q\Leftrightarrow$
+$\{y\in\mathbb{R}^{+}\mid y\log_{a}(b)\in A\}\in q\Leftrightarrow A\in\left(\log_{a}(b)\right)q.$
+Hence $p=\left(\log_{a}(b)\right)q$, as desired.
+$\Leftarrow$)⇐ ) Just observe that $b^{q}=b^{\left(\log_{a}(b)\right)p}$, and $b^{\left(\log_{a}(b)\right)p}=E_{1}\left(b^{\left(\log_{a}(b)\right)},p\right)$ by $(1)$.
+We conclude as $b^{\left(\log_{a}(b)\right)}=a$.
+$(3)$ $\Rightarrow$)⇒ ) This time, consider the function $f(x):=\sqrt[a]{x}:\mathbb{R}^{+}\rightarrow\mathbb{R}$.
+Again, a straightforward computation shows that $\widehat{f}\left(p^{a}\right)=p$, hence $p=\widehat{f}\left(q^{b}\right)$.
+For any $A\subseteq\mathbb{R}^{+}$ we have that
+$A\in\widehat{f}\left(q^{b}\right)\Leftrightarrow(f)^{-1}(A)\in\left(q^{b}\right)\Leftrightarrow\{x\in\mathbb{R}^{+}\mid\{y\in\mathbb{R}^{+}\mid y^{x}\in f^{-1}(A)\}\in q\}\in b\Leftrightarrow$
+$\{y\in\mathbb{R}^{+}\mid y^{b}\in f^{-1}(A)\}\in q\Leftrightarrow\{x\in\mathbb{R}^{+}\mid f\left(x^{b}\right)\in A\}\in q\Leftrightarrow$
+$\{x\in\mathbb{R}^{+}\mid x^{\frac{b}{a}}\in A\}\in q\Leftrightarrow A\in q^{\frac{b}{a}}.$
+$\Leftarrow$)⇐ ) By $(1)$ we have that $q^{b}=E_{1}\left(p^{\frac{a}{b}},b\right)=E_{1}\left(p,\frac{a}{b}\cdot b\right)=p^{a}$.
+$(4)$ Let $A\subseteq\mathbb{R}^{+}$. Then
+$A\in a^{p}\odot a^{q}\Leftrightarrow\{x\in\mathbb{R}^{+}\mid\{y\in\mathbb{R}^{+}\mid xy\in A\}\in a^{q}\}\in a^{p}\Leftrightarrow$
+$\{x\in\mathbb{R}^{+}\mid\{t\in\mathbb{R}^{+}\mid a^{t}\in\{y\in\mathbb{R}^{+}\mid xy\in A\}\}\in q\}\in a^{p}\Leftrightarrow$
+$\{x\in\mathbb{R}^{+}\mid\{t\in\mathbb{R}^{+}\mid x\cdot a^{t}\in A\}\in q\}\in a^{p}\Leftrightarrow\{z\in\mathbb{R}^{+}\mid a^{z}\in\{x\in\mathbb{R}^{+}\mid\{t\in\mathbb{R}^{+}\mid x\cdot a^{t}\in A\}\in q\}\in p\Leftrightarrow$
+$\{z\in\mathbb{R}^{+}\mid\{t\in\mathbb{R}^{+}\mid a^{t+z}\in A\}\in q\}\in p\Leftrightarrow A\in a^{p\oplus q}.$
+$(5)$ Let $A\subseteq\mathbb{R}^{+}$. Then
+$A\in E_{1}(p,q\odot r)\Leftrightarrow\{x\in\mathbb{R}^{+}\mid\{y\in\mathbb{R}^{+}\mid x^{y}\in A\}\in q\odot r\}\in p\Leftrightarrow$
+$\{x\in\mathbb{R}^{+}\mid\{t\in\mathbb{R}^{+}\mid\{z\in\mathbb{R}^{+}\mid x^{tz}\in A\}\in r\}\in q\}\in p\Leftrightarrow$
+$\{x\in\mathbb{R}^{+}\mid\{t\in\mathbb{R}^{+}\mid\{z\in\mathbb{R}^{+}\mid\left(x^{t}\right)^{z}\in A\}\in r\}\in q\}\in p\Leftrightarrow$
+$\{y\in\mathbb{R}^{+}\mid\{z\in\mathbb{R}^{+}\mid s^{z}\in A\}\in r\}\in E_{1}\left(p,q\right)\Leftrightarrow A\in E_{1}\left(E_{1}\left(p,q\right),r\right).$
+$(6)$ This is analogous to $(5)$. Let $A\subseteq\mathbb{R}^{+}$. Then
+$A\in E_{2}\left(p,E_{2}(q,r)\right)\Leftrightarrow\{x\in\mathbb{R}^{+}\mid\{y\in\mathbb{R}^{+}\mid y^{x}\in A\}\in E_{2}(q,r)\}\in p\Leftrightarrow$
+$\{x\in\mathbb{R}^{+}\mid\{t\in\mathbb{R}^{+}\mid\{z\in\mathbb{R}^{+}\mid\left(z^{t}\right)^{x}\in A\}\in r\}\in q\}\in p\Leftrightarrow$
+$\{x\in\mathbb{R}^{+}\mid\{t\in\mathbb{R}^{+}\mid\{z\in\mathbb{R}^{+}\mid z^{tx}\in A\}\in r\}\in q\}\in p\Leftrightarrow\{s\in\mathbb{R}^{+}\mid\{z\in\mathbb{R}^{+}\mid z^{s}\in A\}\in r\}\in p\odot q.\qed$
+Corollary 2.11.
+Let $p,q,\beta\mathbb{N}\setminus{1}$, $a,b\in\mathbb{N}\setminus\{1\}$. Then
+  1. 1.
+    $a^{p}=a^{q}\Leftrightarrow p=q$;
+  2. 2.
+    $p^{a}=p^{b}\Leftrightarrow a=b$.
+Proof.
+(1) This follows from Proposition 2.10.(2).
+(2) This follows from Proposition 2.10.(3), as $p^{\frac{a}{b}}=p$ if and only if $\frac{a}{b}=1$. ∎
+Example 2.12.
+Let $n\in\mathbb{N},n>1$ and $p,q\in\beta\mathbb{N}\setminus\{0\}$ be such that $p\odot q=q\odot p$. Then by Proposition 2.10.we have that
+$E_{1}\left(n^{p},q\right)=n^{p\odot q}=n^{q\odot p}=E_{2}\left(q,n^{p}\right).$
+In particular, this shows the existence of infinitely many pairs of nonprincipal ultrafilters solving $E_{1}(p,q)=E_{2}(q,p)$ (just take, e.g., $p=2^{q}$).
+Proposition 2.10 might again suggest that many arithmetic properties of the exponentiations of natural numbers hold also for $E_{1},E_{2}$.
+For example, Proposition 2.10.(4) generalizes the usual rule $a^{b}\cdot a^{c}=a^{bc}$ to the case where $b,c$ are ultrafilters.
+However, this same rule cannot be generalized to the case where $a$ is an ultrafilter, as we show in the next proposition.
+Proposition 2.13.
+Let $p\in\beta\mathbb{N}\setminus\{\mathbb{N}\}$, and let $a,b\in\mathbb{N}$. Then $p^{a}\odot p^{b}\neq p^{a+b}$.
+Proof.
+Assume, by contrast, that $p^{a}\odot p^{b}=p^{a+b}$.
+Let $F:\mathbb{N}\rightarrow\mathbb{N}$ be the function that maps a number $n$ to its largest prime divisor.
+Observe that, for all $p_{1},p_{2}\in\beta\mathbb{N}$, $\hat{\Omega}(E_{2}(p_{1},p_{2}))=p_{1}\odot\hat{\Omega}(p_{2})$ and $\hat{\Omega}(p_{1}\odot p_{2})=\hat{\Omega}(p_{1})\oplus\hat{\Omega}(p_{2})$.
+We consider two cases.
+Case 1: $\hat{F}(p)=i\in\mathbb{N}$.
+In this case, consider the function $\Omega:\mathbb{N}\rightarrow\mathbb{N}$ that maps a number $n$ to the sum of the exponents in its prime factorization.
+Necessarily, $\hat{\Omega}(p)\notin\mathbb{N}$; in fact, if by contrast $\hat{\Omega}(p)=j\in\mathbb{N}$, we would have that $p\in\overline{S}$, where
+$S=\{n\in\mathbb{N}\mid F(n)=i,\Omega(n)=j\},$
+which is a contradiction as $p$ is nonprincipal and $S$ is finite. Hence $\hat{\Omega}(p)\notin\mathbb{N}$. As $p^{a}\odot p^{b}=p^{a+b}$, we have that
+$\hat{\Omega}\left(p^{a}\odot p^{b}\right)=\hat{\Omega}\left(p^{a}\right)\oplus\hat{\Omega}\left(p^{b}\right)=a\hat{\Omega}(p)\oplus b\hat{\Omega},$
+whilst
+$\hat{\Omega}\left(p^{a+b}\right)=(a+b)\hat{\Omega}(p).$
+We conclude as $a\hat{\Omega}(p)\oplus b\hat{\Omega}(p)\neq(a+b)\hat{\Omega}(p)$ by Theorem 2.6.
+Case 2: $\hat{F}(p)\notin\mathbb{N}$.
+Let $G:\mathbb{N}\rightarrow\mathbb{N}$ be the function $n\rightarrow\max\{m\in\mathbb{N}\mid F(n)^{m}\ \text{divides}\ n\}$, and let $H(n):=F(n)^{G(n)}$.
+As $\hat{F}(p)$ is nonprincipal, also $\hat{H}(p)$ is nonprincipal.
+And
+$\hat{H}\left(p^{a}\odot p^{b}\right)=b\hat{H}(p)\neq(a+b)\hat{H}(p)=\hat{H}\left(p^{a+b}\right),$
+again by Theorem 2.6. ∎
+Remark 2.14.
+The above results prove that the equalities $E_{1}(r,p)\odot E_{1}(r,q)=E_{1}(r,p\oplus q),E_{2}(p,r)\odot E_{2}(q,r)=E_{2}(p\oplus q,r)$ does not hold when $r$ is principal.
+They do not need to hold even when $r$ is nonprincipal.
+E.g., assume that $r$ is nonprincipal and $P=\{n\in\mathbb{N}\mid n\ \text{is a prime}\}\in r$.
+A direct computation shows that
 $I_{1}=\{n\in\mathbb{N}\mid\exists p_{1},p_{2}\in P,a,b\in\mathbb{N}\ \text{s.t.}\ p_{1}<a<p_{2}<b\ \text{and}\ p_{1}^{a}p_{2}^{b}=n\}\in E_{1}(r,p)\odot E_{1}(r,q),$
@@
 $J=\{n\in\mathbb{N}\mid\exists p\in P,a\in\mathbb{N}\ \text{s.t.}\ p^{a}=n\}\in E_{1}(r,p\oplus q)\cap E_{2}(p\oplus q,r),$
-and we conclude as $I_{1}\cap J=I_{2}\cap J=\emptyset.$
-$(5)$ Let $A\subseteq\mathbb{N}$. Then
-$A\in E_{1}(p,q\odot r)\Leftrightarrow\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid n^{m}\in A\}\in q\odot r\}\in p\Leftrightarrow$
-$\{n\in\mathbb{N}\mid\{a\in\mathbb{N}\mid\{b\in\mathbb{N}\mid n^{ab}\in A\}\in r\}\in q\}\in p\Leftrightarrow\{n\in\mathbb{N}\mid\{a\in\mathbb{N}\mid\{b\in\mathbb{N}\mid\left(n^{a}\right)^{b}\in A\}\in r\}\in q\}\in p\Leftrightarrow$
-$\{n\in\mathbb{N}\mid\{a\in\mathbb{N}\mid\{b\in\mathbb{N}\mid\left(n^{a}\right)^{b}\in A\}\in r\}\in q\}\in p\Leftrightarrow\{c\in\mathbb{N}\mid\{d\in\mathbb{N}\mid c^{d}\in A\}\in r\}\in E_{1}\left(p,q\right)\Leftrightarrow$
-$A\in E_{1}\left(E_{1}\left(p,q\right),r\right).$
-$(6)$ This is analogous to $(5)$. Let $A\subseteq\mathbb{N}$. Then
-$A\in E_{2}\left(p,E_{2}(q,r)\right)\Leftrightarrow\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid m^{n}\in A\}\in E_{2}(q,r)\}\in p\Leftrightarrow$
-$\{n\in\mathbb{N}\mid\{a\in\mathbb{N}\mid\{b\in\mathbb{N}\mid\left(b^{a}\right)^{n}\in A\}\in r\}\in q\}\in p\Leftrightarrow\{n\in\mathbb{N}\mid\{a\in\mathbb{N}\mid\{b\in\mathbb{N}\mid b^{an}\in A\}\in r\}\in q\}\in p\Leftrightarrow$
-$\{c\in\mathbb{N}\mid\{b\in\mathbb{N}\mid b^{c}\in A\}\in r\}\in p\odot q.\qed$
-Example 2.5.
-Let $n\in\mathbb{N},n>1$ and $p,q\in\beta\mathbb{N}\setminus\{0\}$ be such that $p\odot q=q\odot p$. Then by Proposition 2.4.we have that
-$E_{1}\left(n^{p},q\right)=n^{p\odot q}=n^{q\odot p}=E_{2}\left(q,n^{p}\right).$
-In particular, this shows the existence of infinitely many pairs of nonprincipal ultrafilters solving $E_{1}(p,q)=E_{2}(q,p)$ (just take, e.g., $p=2^{q}$).
-Their definition should make evident that $E_{1},E_{2}$ are ultrafilters operations with relationships with exponential configurations.
+and we conclude that $E_{1}(r,p)\odot E_{1}(r,q)\neq E_{1}(r,p\oplus q)$ and $E_{2}(p,r)\odot E_{2}(q,r)\neq E_{2}(p\oplus q,r)$ as $I_{1}\cap J=I_{2}\cap J=\emptyset.$
+Their definition should make evident that $E_{1},E_{2}$ are ultrafilters operations that can be used to study the partition regularity of certain exponential configurations.
 Example 2.15.
@@
 $B_{0}=\{m\in\mathbb{N}\mid\{n\in\mathbb{N}\mid\{m^{n}\in A\}\in p\}\in p.$
-Now take $F_{1}\in\mathcal{P}_{f}(B_{0})$ and let
+For $F_{1}\in\mathcal{P}_{f}(B_{0})$, let
 $B_{1}=\bigcap_{t\in F_{1}}\{n\in\mathbb{N}\mid t^{n}\in A\}\cap B_{0}\in p.$
-Proceeding inductively like above, we construct a sequence $\langle F_{i}\rangle_{i=1}^{\infty}$ of finite subsets of $B_{0}$ with the property that for all $j<i$ one has that
-$F_{j}^{F_{i}}=\{m^{n}\mid m\in F_{j},n\in F_{i}\}\in A.$
+Proceeding inductively like above, it is possible to construct a sequence $\langle F_{i}\rangle_{i=1}^{\infty}$ of finite subsets of $B_{0}$ such that, for all $j<i$, $F_{j}^{F_{i}}=\{m^{n}\mid m\in F_{j},n\in F_{i}\}\in A$.
 For example, if $F_{i}=\{a_{i}\}$ for all $i$, we obtain an infinite subset $X=\{a_{i}\mid i\in\mathbb{N}\}$ of $B_{0}$ such that $\{a_{j}^{a_{i}}\mid a_{i},a_{j}\in X,j<i\}\subseteq A$.
-Moreover, by choosing $p$ with additional properties, one could get more complicated configurations.
-For example, by taking $p$ such that each set in $p$ contains arbitrarily long arithmetic progressions, we can take all sets $F_{i}$ to be arithmetic progressions of length $i$ in $B$.
-Example 2.7.
-If we argue as above with $E_{2}$ instead of $E_{1}$, we get analogous results with the order relationship between $i$ and $j$ switched.
-So, in the same notations, we get a sequence $\langle F_{i}\rangle_{i=1}^{\infty}$ of finite subsets of $B_{0}$ with the property that for all $j<i$ one has that
-$F_{j}^{F_{i}}=\{m^{n}\mid m\in F_{j},n\in F_{i}\}\in A.$
-For example, if $F_{i}=\{a_{i}\}$ for all $i$, we obtain an infinite subset $X=\{a_{i}\mid i\in\mathbb{N}\}$ of $B_{0}$ such that $\{a_{j}^{a_{i}}\mid a_{i},a_{j}\in X,j>i\}\subseteq A$.
+Arguing as above with $E_{2}$ instead of $E_{1}$, we could deduce an analogous results with the order relationship between $i$ and $j$ switched.
 The reader with a basic knowledge of ultrafilters will recognize that a trivial way to strengthen the above examples would be to take $p$ such that $E_{1}(p,p)=p$ or $E_{2}(p,p)=p$, namely some kind of exponential idempotent.
 We devote Section 2.3 to the study of the existence of such ultrafilters.
-To conclude this section, we want to show that the operations $E_{1},E_{2}$ are related to exponential triples.
-This should not come as a surprise as, even if using a different language, Di Nasso and Ragosta in [2] to prove the exponential Schur theorem employed ultrafilters $u$ of the form $u=2^{q}$ where $q$ had the form $q=2^{p}\odot p$ for some $p$ witness of van der Waerden Theorem.
-In [4], the authors proved the following result, which allows to generalize the observation of Di Nasso and Ragosta::
-Theorem 2.8.
-Let $A$ be a multiplicatively piecewise syndetic set. Then $A$ contains an exponential triple.
-As an immediate corollary, the following holds.
-Corollary 2.9.
-Let $\mathcal{U}\in\overline{K(\beta\mathbb{N},\odot)}$. Then every $A\in\mathcal{U}$ contains an exponential triple.
-Notice that, by Corollary 2.9, it is simple to prove that several configurations are partition regular by using [1, Lemma 2.1].
-To recall it, we need to introduce the following notation: for $p\in\beta\mathbb{N}$, $f\left(x_{1},\dots,x_{n}\right):\mathbb{N}^{n}\rightarrow\mathbb{N}$, we write $p\models f\left(x_{1},\dots,x_{n}\right)=0$ if for all $A\in\mathcal{U}$ there exists $a_{1},\dots,a_{n}\in A$ such that $f\left(a_{1},\dots,a_{n}\right)=0$.
-Lemma 2.10.
-[1, Lemma 2.1] Let $p\in\beta\mathbb{N}$, and let $f\left(x_{1},\dots,x_{n}\right),g\left(y_{1},\dots,y_{m}\right)$ be two functions such that $p\models\left(f_{1}\left(x_{1},\dots,x_{n}\right)=0\right)\wedge\left(g\left(y_{1},\dots,y_{m}\right)=0\right)$.
-Then $p\models\left(f_{1}\left(x_{1},\dots,x_{n}\right)=0\right)\wedge\left(g\left(y_{1},\dots,y_{m}\right)=0\right)\wedge\left(x_{1}=y_{1}\right)$.
-As said, the above result allows us to recover the partition regularity of many equations and configurations. In fact, pick $p\in\overline{K(\beta\mathbb{N},\odot)}$ multiplicatively idempotent. Then:
-[…]
-  • as $p$ is idempotent, $p\models y_{1}\dot{}\dots\dot{y}_{m}=z$;
+We now want to show that the operations $E_{1},E_{2}$ are related to the partition regularity of exponential triples, as well as of other exponential configurations.
+This should not come as a surprise: the main results in [2] can be restated in terminology introduced in Section 2.1 as follows:
+Theorem 2.16 ([2, Theorem 2.1, restated]).
+If $p$ is a Van der Waerden’s witness then $2^{p}\odot p\models\{x,y,2^{x}y\}$.
+From this fact, in [2, Corollary 2.3] the authors deduced the exponential Schur’s Theorem by the $\log_{2}$-partition method mentioned in the introduction.
+Their proof actually shows the following fact:
+Theorem 2.17 ([2, Theorem 2.3, restated]).
+If $p$ is Van der Waerden’s witness then $2^{2^{p}\odot p}\models\{x,y,x^{y}\}$.
+As it is well-known (see e.g.
+[5, Theorem 14.5]), any $p\in\overline{K(\beta\mathbb{N},\odot)}$ is a Van der Waerden’s witness; moreover, as $p\in\overline{K(\beta\mathbb{N},\odot)}$, we have that $2^{p}\odot p\in\overline{K(\beta\mathbb{N},\odot)}$.
+This simple observation, joint with Theorem 2.17 and Lemma 2.4, allows to prove easily that several configurations are partition regular.
+In fact, let $p\in\overline{K(\beta\mathbb{N},\odot)}$, and let $q=2^{p}$.
+Then:
 […]
 […]
-and so on, as we can always substitute a variable $x$ with a polynomial $P\left(y_{1},\dots,y_{n}\right)$ whenever $p\models x=P\left(y_{1},\dots,y_{n}\right)$.
+  • by furtherly assuming that $p=p\odot p$, one gets that $q\models x^{\log_{2}(y)}=z$, hence e.g. $q\models x_{1}^{x_{2}\log_{2}x_{3}}=y_{1}\cdot y_{2}^{y_{3}}$;
+  • if $p$ is also combinatorially rich, then $p\models x=y+z^{2}$ (see e.g. [1, Theorem 2.11], hence $q\models x=yt^{\log_{2}(t)}$,
+and so on, as we can always substitute a variable $x$ with a polynomial $P\left(y_{1},\ldots,y_{n}\right)$ whenever $q\models x=P\left(y_{1},\ldots,y_{n}\right)$.
 We refer to [1, Theorem 2.11] for a long list of such admissible substitutions.
-Corollary 2.9 can be generalized also in another direction. Let us first fix a definition.
-Definition 2.11.
-We let
-$EVdW=\{p\in\beta\mathbb{N}\mid\forall A\in p,\forall k\in\mathbb{N}\,\exists\,x,y\in A\ \text{s.t.}\ x^{y},\dots,x^{\left(y^{k}\right)}\in A\}.$
+To conclude this section we study some properties of the set of exponential Schur’s witnesses, namely
+$\mathrm{ESW}=\{p\in\beta\mathbb{N}\mid\forall A\in p\ \exists x,y\in A\ x^{y}\in A\}.$
+The set $\mathrm{ESW}$ has some nice closure property.
+To prove them, we will need the following strengthening of Di Nasso and Ragosta’s results, whose proofs are almost identical to those of the results in [2].
+We add them here for completeness.
 Theorem 2.18.
-For all $\mathcal{U}\in\overline{K(\beta\mathbb{N},\odot)}$, for all $n\in\mathbb{N}_{>1}$, $n^{\mathcal{U}}\in EVdW$.
-Actually, Theorem 2.12 can be seen as a very particular case of a much more general phenomenon involving $E_{1},E_{2}$.
-Theorem 2.13.
-Let $p,q\in\beta\mathbb{N}\setminus\{0\}$. Then:
-  1. 1.
-    If $q\in\overline{K(\beta\mathbb{N},\odot)}$ then $E_{1}\left(p,q\right)\in EVdW$. In particular, $E_{1}\left(p,q\right)$ witnesses the partition regularity of exponential triples;
-  2. 2.
-    If $q\in EVdW$ then $E_{2}(p,q)\in EVdW$. In particular, $E_{2}(p,q)$ witnesses the partition regularity of exponential triples.
+Let $n\in\mathbb{N}$. If $p$ is a Van der Waerden’s witness then $n^{p}\odot p\models\{x,y,n^{x}y\}$.
 Proof.
-(1) Let $k\in\mathbb{N}$ and let $A\in E_{1}(p,q)$. We will show that $A$ contains a configuration of the form $x,y,x^{y},\dots,x^{\left(y^{k}\right)}$.
-By definition, $I=\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid n^{m}\in A\}\in q\}\in p$.
+Let $A\in n^{p}\odot p$.
+By definition, this means that $A^{\prime}=\{a\in\mathbb{N}\mid A/n^{a}\in p\}\in p$, where $A/n^{a}=\{m\in\mathbb{N}\mid n^{a}m\in A\}$.
+Pick $a\in A^{\prime}$.
+As $A^{\prime}\cap A/n^{a}\in p$, there are $b,c,b+n^{a}c\in A/n^{a}\cap A^{\prime}$.
+In particular, $A/n^{b},A/n^{b+n^{a}c}\in p$.
+Let $d\in A/n^{b}\cap A/{n^{b+n^{a}c}}$.
+Letting $x=n^{a}c,y=n^{b}d$, we have that $x\in A$ as $c\in A/n^{a}$, $y\in A$ as $d\in A/n^{b}$, and $n^{x}y=n^{n^{a}c}n^{b}d=n^{b+n^{a}c}d\in A$ as $d\in A/n^{b+n^{a}c}$.∎
+Corollary 2.19.
+Let $n\in\mathbb{N}$. If $p$ is a Van der Waerden’s witness then $n^{n^{p}\odot p}\in\mathrm{ESW}$.
+Proof.
+Let $A\in n^{n^{p}\odot p}$.
+Let $I_{A}\in\{m\mid n^{m}\in A\}\in n^{p}\odot p$.
+Let $a,b\in I_{A}$ be such that $n^{a}b\in I_{A}$.
+Then $x:=n^{b},y:=n^{a}$ are elements of $A$ such that $x^{y}=\left(n^{b}\right)^{\left(n^{a}\right)}=n^{bn^{a}}\in A$.
+∎
+Our first result is that Corollary 2.19 can be easily generalized as follows.
+Proposition 2.20.
+Let $q\in\beta\mathbb{N}\setminus\{0,1\}$. If $p$ is a Van der Waerden’s witness then $E_{1}\left(q,E_{1}(q,p)\odot p\right)\in\mathrm{ESW}$.
+Proof.
+Let $A\in E_{1}\left(q,E_{1}(q,p)\odot p\right)$. By definition,
+$\underbrace{\{n\in\mathbb{N}\mid\overbrace{\{m\in\mathbb{N}\mid\{a\in\mathbb{N}\mid\{b\in\mathbb{N}\mid n^{m^{a}b}\in A\}\in p\}\in p\}}^{J}\in q\}}_{I}\in q.$
+If $n\in I\cap J$ then $A\in n^{n^{p}\odot p}$, and we conclude by Corollary 2.19.∎
+Similarly, a more restrictive condition on $p$ allows to prove the following:
+Proposition 2.21.
+Let $p,q\in\beta\mathbb{N}\setminus\{0,1\}$. Assume that, for all $n>1$, $p\models\{x,y,xn^{y}\}$. Then $E_{1}\left(q,p\right)\in\mathrm{ESW}$.
+Proof.
+Let $A\in E_{1}(q,p)$.
+By definition, $I=\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid n^{m}\in A\}\in p\}\in q$.
 Let $n\in I$ and $J_{n}=\{m\in\mathbb{N}\mid n^{m}\in A\}$.
-Take $a,b\in J_{n}$ with $a,b,an^{b},\dots,an^{kb}\in J_{n}$, whose existence is ensured by Corollary 2.9 since, as $J_{n}\in q\in\overline{K(\beta\mathbb{N},\odot)}$, $J_{n}$ is multiplicatively piecewise syndetic set.
-Therefore,
-$n^{a},n^{b},n^{an^{b}}=\left(n^{a}\right)^{\left(n^{b}\right)},\dots,n^{an^{kb}}=\left(n^{a}\right)^{\left(\left(n^{b}\right)^{k}\right)}\in A,$
-and we conclude by letting $x=n^{a},y=n^{b}$.
-(2) Let $q\in EVdW$, let $k\in\mathbb{N}$ be given and let $A\in E_{2}(p,q)$.
-By definition, $I=\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid m^{n}\in A\}\in q\}\in p$.
-Let $n\in I$ and let $J_{n}=\{m\in\mathbb{N}\mid m^{n}\in A\}$.
-Take $a,b\in J_{n}$ with $a,b,a^{\left(b^{n}\right)},\dots,a^{\left(b^{kn}\right)}\in J_{n}$.
-Hence
-$a^{n},b^{n},\left(a^{\left(b^{n}\right)}\right)^{n}=a^{nb^{n}}=\left(a^{n}\right)^{\left(b^{n}\right)},\dots,\left(a^{\left(b^{kn}\right)}\right)^{n}=a^{nb^{kn}}=\left(a^{n}\right)^{\left(\left(b^{n}\right)^{k}\right)}\in A,$
-and we conclude by letting $x=a^{n},y=b^{n}$.∎
-As a consequence of [6, Theorem 6], if $Sc_{+}=\{p\in\beta\mathbb{N}\mid\forall A\in p\ \exists a,b\in A\ a+b\in A\}$, then $Sc_{+}\cap EVdW=\emptyset$; in particular,
-$\overline{K(\beta\mathbb{N},\odot)}\cap EVdW=\emptyset,\overline{E(\oplus)}\cap EVdW=\emptyset,$
+Take $a,b\in J_{n}$ with $a,b,an^{b}\in J_{n}$.
+Therefore
+$n^{a},n^{b},n^{an^{b}}=\left(n^{a}\right)^{\left(n^{b}\right)}\in A,$
+and we conclude by letting $x=n^{a},y=n^{b}$. ∎
+By [5, Theorem 4.40], any $q\in\overline{K(\beta\mathbb{N},\odot)}$ consists solely of multiplicatively piecewise syndetic sets; hence, by Theorem 2.5, any such $q$ satisfies that, for all $n>1$, $q\models\{x,y,xn^{y}\}$.
+Therefore, Theorem 2.21 has the following immediate corollary:
+Corollary 2.22.
+Let $q\in\overline{K(\beta\mathbb{N},\odot)}$. Then for all $p\in\beta\mathbb{N}\setminus\{0,1\}$ we have that $E_{1}\left(p,q\right)\in\mathrm{ESW}$.
+Remark 2.23.
+If $Sc_{+}=\{p\in\beta\mathbb{N}\mid\forall A\in p\ \exists a,b\in A\ a+b\in A\}$, Theorem 2.7 tells us that $Sc_{+}\cap\mathrm{ESW}=\emptyset$; in particular, if $E(\oplus)=\{p\in\beta\mathbb{N}\mid p\oplus p=p\}$ then
+$\overline{K(\beta\mathbb{N},\odot)}\cap\mathrm{ESW}=\emptyset,\overline{E(\oplus)}\cap\mathrm{ESW}=\emptyset,$
 as $\overline{K(\beta\mathbb{N},\odot)}\cup\overline{E(\oplus)}\subseteq Sc_{+}$.
-Then, in particular, by Theorem 2.12 we also deduce that for all $n>1$, for all $p\in\overline{K(\beta\mathbb{N},\odot)}$, $n^{p}\notin\overline{K(\beta\mathbb{N},\odot)}$.
+By Theorem 2.21.(1), we get that if $q\in\overline{K(\beta\mathbb{N},\odot)}$ then $E_{1}\left(p,q\right)\notin\overline{K(\beta\mathbb{N},\odot)}$; in particular, for $p,q\in\overline{K(\beta\mathbb{N},\odot)}$, $E_{1}(p,q)\notin\overline{K(\beta\mathbb{N},\odot)}$: the smallest bilateral multiplicative ideal of $\beta\mathbb{N}$ is very much not closed under the $E_{1}$ exponentiation.
 Non existence of exponential idempotents
-As it is nowadays well-known, idempotents are often key ingredients in the search for infinite partition regular structures, the most famous example being the relationship between idempotent ultrafilters with respect to the sum and Hindman Theorem (see [5, Chapter 5]).
-As Examples 2.6, 2.7 suggest, having introduced the binary operations $E_{1},E_{2}$, it makes sense to study if they admit idempotents:
+Having introduced two natural binary operations on $\beta\mathbb{N}$, it makes sense to ask if they admit idempotents, defined as follows.
 Definition 2.24.
 We say that $p\in\beta\mathbb{N}$ is $E_{1}$-idempotent (resp. $E_{2}$-idempotent) if $E_{1}(p,p)=p$ (resp. if $E_{2}(p,p)=p$).
-First of all, since $E_{1},E_{2}$ are not associative, we cannot approach the existence of idempotents problem via the usual straightforward application of Ellis’ Theorem (see [5, Theorem 2.5.] for details).
+First of all, since $E_{1},E_{2}$ are not associative, we cannot approach the existence of idempotents problem via the usual method of applying Ellis’ Theorem (see [5, Theorem 2.5] for details).
 Moreover, since $1$ is the only principal $E_{1}-$idempotent, and the only $E_{2}$-idempotent as well, from now on we will use “idempotent” to actually mean “nonprincipal idempotent”.
-In the following we want to study properties of generalizations of the operation $(p,q)\rightarrow 2^{p}\odot q$, first introduced by Di Nasso and Ragosta in [2].
-We will use the following notations: for $n\in\mathbb{N},p,q\in\beta\mathbb{N}$ we let:
-$p\oleft_{n}q:=n^{p}\odot q;p\oright_{n}q:=p\odot n^{q}.$
-The following result shows that the existence of idempotents for $E_{2}$ is related to that of ⊣⃝$\oleft_{n}⊣$-idempotents, whilst the existence of idempotents for $E_{1}$ is related to that of ⊢⃝$\oright_{n}⊢$-idempotents.
-Proposition 2.15.
-If $n\in\mathbb{N}_{>1}$, then $(1)\Rightarrow(2)\Rightarrow(3)$, where
-  1. (i)
-    there exists $p\in\beta\mathbb{N}$ such that $p\oplus n^{p}=p$;
-  2. (ii)
-    there exists $p\in\beta\mathbb{N}$ such that $p\oright_{n}p=p$;
-  3. (iii)
-    there exists $p\in\beta\mathbb{N}$ such that $E_{1}(p,p)=p$.
-Analogously, $(4)\Rightarrow(5)\Rightarrow(6)$, where
-  1. (iv)
-    there exists $p\in\beta\mathbb{N}$ such that $n^{p}\oplus p=p$;
-  2. (v)
-    there exists $p\in\beta\mathbb{N}$ such that $p\oleft_{n}p=p$;
-  3. (vi)
-    there exists $p\in\beta\mathbb{N}$ such that $E_{2}(p,p)=p$.
-Proof.
-All the implications follows by taking $q$ as given by the hypothesis, by letting $p=n^{q}$ and by applying Proposition 2.4.
-For this reason, we prove only that $(v)$ entails $(vi)$, as the other implications are very similar.
-$(v)\Rightarrow(vi)$ Let $q\in\beta\mathbb{N}$ be such that $q\oleft_{n}q=q$ and let $p=n^{q}$.
-Then $E_{2}(p,p)=E_{2}\left(n^{q},n^{q}\right)$.
-As $n^{q}=E_{2}(q,n)$, by Proposition 2.4.(6) we have that
-$E_{2}\left(n^{q},n^{q}\right)=E_{2}\left(n^{q},E_{2}\left(q,n\right)\right)=E_{2}\left(n^{q}\odot q,n\right)=E_{2}\left(q,n\right)=p.\qed$
-Moreover, in the search of $E_{1}$- and $E_{2}$-idempotent ultrafilters, we make the following observation:
+Our starting observation is the following:
 Proposition 2.25.
@@
 Proof.
-We just have to apply Proposition 2.4.(5)-(6).
-In fact, if $p=p\odot E_{1}(q,p)$ then
-$E_{1}(q,p)=E_{1}(q,p)=E_{1}(q,p\odot E_{1}(q,p))=E_{1}\left(E_{1}(q,p),E_{1}(q,p)\right)$
+We just have to apply Proposition 2.10.(5)-(6). In fact, if $p=p\odot E_{1}(q,p)$ then
+$E_{1}(q,p)=E_{1}(q,p\odot E_{1}(q,p))=E_{1}\left(E_{1}(q,p),E_{1}(q,p)\right)$
 and, if $p=E_{2}(p,q)\odot p$ then
 $E_{2}(p,q)=E_{2}\left(E_{2}(p,q)\odot p,q\right)=E_{2}\left(E_{2}(p,q),E_{2}(p,q)\right).\qed$
-All that has been said so far might point to the actual existence of exponential idempotents.
-Actually, as we are going to show, this is not the case for $E_{1}$, and at least not trivially the case for $E_{2}$.
+In particular, the existence of $n\in\mathbb{N}_{>1}$ and $p\in\beta\mathbb{N}\setminus\mathbb{N}$ with $p\odot n^{p}=p$ would force $n^{p}$ to be an $E_{1}$-idempotent; similarly, if $n^{p}\odot p=p$ then $n^{p}$ would be $E_{2}$-idempotent.
+However, idempotents cannot exist for $E_{1}$, and cannot exist for $E_{2}$ under certain additional assumptions, as we are now going to show.
 As for $E_{1}$, the non existence of idempotents follows from an even stronger fact, that we now prove.
@@
 Proof.
-By [6, Lemma 23], there is a $4$-coloring $\mathbb{N}=A_{1}\cup\dots\cup A_{4}$ of $\mathbb{N}$ such that there is no monochromatic pair $b,a^{b}$ with $\log_{2}a\leq b$ and $a,b\geq 1$.
+By [7, Lemma 23], there is a $4$-coloring $\mathbb{N}=A_{1}\cup\ldots\cup A_{4}$ of $\mathbb{N}$ such that there is no monochromatic pair $b,a^{b}$ with $\log_{2}a\leq b$ and $a,b\geq 1$.
 Now assume that $E_{1}(p,q)=q$; by definition, this means that for all $A\subseteq\mathbb{N}$ we have that
 $A\in q\Leftrightarrow\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid n^{m}\in A\}\in q\}\in p.$
-Now take $A_{i}\in\{A_{1},\dots,A_{4}\}$ such that $A_{i}\in q$.
+Now take $A_{i}\in\{A_{1},\ldots,A_{4}\}$ such that $A_{i}\in q$.
 In particular, $B=\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid n^{m}\in A_{i}\}\in q\}\neq\emptyset$, so we can pick $n\in B$.
@@
 As $q$ is non principal, $I_{n}$ must contain some $m>\log_{2}(n)$.
-But $m,n^{m}\in A_{i}$ with $\log_{2}(n)\leq m$, which is absurd.
+But $m,n^{m}\in A_{i}$ with $\log_{2}(n)\leq m$, which is a contradiction.
 ∎
-Putting together Theorem 2.17 and Propositions 2.15, 2.16, we get the following:
+Putting together Theorem 2.26 and Proposition 2.25, we get the following:
 Corollary 2.27.
-For all $n>1\in\mathbb{N}$ and $p,q\in\beta\mathbb{N}\setminus\{0,1\}$ we have that $p\oplus n^{p}\neq p$ and $p\odot E_{1}(q,p)\neq p$.
-Notice that, in particular, also $p\oright_{n}{p}=p$ has no solution in $\beta\mathbb{N}$ if $n>1$.
+For all $n>1\in\mathbb{N}$ and $p,q\in\beta\mathbb{N}\setminus\{0,1\}$ we have that $p\odot n^{p}\neq p$ and $p\odot E_{1}(q,p)\neq p$.
 Naively, from Theorem 2.26 it would be natural to imagine that some similar argument could be used to prove the following conjecture:
@@
 By contrast, let $p$ be such an ultrafilter.
-Let $S:\mathbb{N}\rightarrow\mathbb{N}$ be the function that maps $n=p_{1}^{a_{1}}\cdot\dots\cdot p_{k}^{a_{k}}$ to $\sum_{i\leq k}a_{i}$, where $n=p_{1}^{a_{1}}\cdot\dots\cdot p_{k}^{a_{k}}$ is the prime decomposition of $n$.
-Observe that $S(E_{2}(p_{1},p_{2}))=p_{1}\odot S(p_{2})$ and $S(p_{1}\odot p_{2})=S(p_{1})\oplus S(p_{2})$ for all $p_{1},p_{2}\in\beta\mathbb{N}$.
-In particular, $S(p)=S(E_{2}(p,p))=p\odot S(p)$ and $S(p)=S(p\odot p)=S(p)\oplus S(p)$, therefore $S(p)$ is additively idempotent and $p\odot S(p)=S(p)\oplus S(p)$.
-If $S(p)\in\mathbb{N}$, then by additive idempotency we would get that $S(p)=0$, from which $p=1$, which is absurd as $p\in\beta\mathbb{N}\setminus\mathbb{N}$.
-So $S(p)$ is additively idempotent and nonprincipal, so $n\mathbb{N}\in S(p)$ for all $n\in\mathbb{N}$.
-Therefore we can apply [5, Theorem 13.14] to deduce that it must be $p\odot S(p)\neq S(p)\oplus S(p)$, reaching the desired contrast.
+Let $\Omega:\mathbb{N}\rightarrow\mathbb{N}$ be the function that maps $n$ to the sum of the exponents in its prime factorization.
+In particular, $\hat{\Omega}(p)=\hat{\Omega}(E_{2}(p,p))=p\odot\hat{\Omega}(p)$ and $\hat{\Omega}(p)=\hat{\Omega}(p\odot p)=\hat{\Omega}(p)\oplus\hat{\Omega}(p)$, therefore $\hat{\Omega}(p)$ is additively idempotent and $p\odot\hat{\Omega}(p)=\hat{\Omega}(p)\oplus\hat{\Omega}(p)$.
+If $\hat{\Omega}(p)\in\mathbb{N}$, then by additive idempotency we would get that $\hat{\Omega}(p)=0$, from which $p=1$, which is a contradiction as $p\in\beta\mathbb{N}\setminus\mathbb{N}$.
+So $\hat{\Omega}(p)$ is additively idempotent and nonprincipal, hence $n\mathbb{N}\in\hat{\Omega}(p)$ for all $n\in\mathbb{N}$.
+Thus we can apply [5, Theorem 13.14] to deduce that $p\odot\hat{\Omega}(p)\neq\hat{\Omega}(p)\oplus\hat{\Omega}(p)$, reaching the desired contradiction.
 ∎
@@
 Question 3.1.
-Let $p\in E\left(K(\beta\mathbb{N},\odot)\right)$ be such that each member of $p$ is also additive central.
+Let $p\in E\left(K(\beta\mathbb{N},\odot)\right)$ be such that each member of $p$ is also additively central.
 Is it true or not that each $A\in p$ contains a configuration of the form $\{m,n,m+n,m2^{n}\}$?
@@
 Question 3.2.
-Are there $p,q\in\beta\mathbb{N}\setminus\mathbb{N}$ that solve any of the following equations
+Prove that there exist $p,q,r\in\beta\mathbb{N}\setminus\mathbb{N}$ that solve any of the following equations
   1. 1.
@@
   6. 6.
+    $E_{1}(r,p)\odot E_{1}(r,q)=E_{1}(r,p\oplus q)$;
+  7. 7.
     $E_{2}(q,p)\odot p=p$;
-  7. 7.
-    $E_{2}(q,p)\odot q=q$.
+  8. 8.
+    $E_{2}(q,p)\odot q=q$;
+  9. 9.
+    $E_{2}(p,r)\odot E_{2}(q,r)=E_{2}(p\oplus q,r)$.
+For one ultrafilter, the following question, for which we conjecture a negative answer, seems relevant:
+Question 3.3.
+Does it exist $p\in\beta\mathbb{N}\setminus\mathbb{N}$ such that $E_{1}(p,p)=E_{2}(p,p)$?
 Of course, as discussed in the paper, more relevant questions are the following:
-Question 3.3.
-Does it exist $p$ that is $E_{2}$-idempotent? Does it exist $p$ that is ⊣⃝$\oleft_{n}⊣$-idempotent?
+Question 3.4.
+Does it exist $p$ that is $E_{2}$-idempotent?
 Finally, as a general direction connected to all the above questions, we believe that it would be interesting to further study the properties of exponentially rich ultrafilters; of course, this rather vague idea could be made precise in several different ways.
 We believe it would be interesting to start by studying the following:
-Question 3.4.
-What are the algebraical properties of $EVdW$, besides Theorem 2.12?
+Question 3.5.
+What are the algebraical properties of $\mathrm{ESW}$?

### v5

@@
 1   Introduction
-A large class of problems in arithmetic Ramsey theory regards the so-called partition regular configurations and equations on =$\mathbb{N}=\{1,2,\ldots\}$.
+A large class of problems in arithmetic Ramsey theory regards the so-called partition regular configurations and equations on =$\mathbb{N}=\{1,2,\ldots\}$. We let $\omega=\mathbb{N}\cup\{0\}$.
 Definition 1.1.
@@
 We say that the configuration $\{f_{1}\left(x_{1},\ldots,x_{n}\right),\ldots,f_{m}\left(x_{1},\ldots,x_{n}\right)\}$ is partition regular on $\mathbb{N}$ (PR from now on) if for all finite partitions $\mathbb{N}=A_{1}\cup\ldots\cup A_{k}$ there are $i\leq k$ and $x_{1},\ldots,x_{n}\in\mathbb{N}$ such that for all $j\leq m$ $f_{j}\left(x_{1},\ldots,x_{n}\right)\in A_{i}$.
-Analogously, we say that the system $\sigma\left(x_{1},\ldots,x_{n}\right)$ defined by
+Analogously, if $f_{1}\left(x_{1},\ldots,x_{n}\right),\ldots,f_{m}\left(x_{1},\ldots,x_{n}\right):\mathbb{N}^{n}\rightarrow\omega$ and the system $\sigma\left(x_{1},\ldots,x_{n}\right)$ is defined by
 $\begin{cases}f_{1}\left(x_{1},\ldots,x_{n}\right)=0,\\ \ \ \ \ \ \ \ \ \ \vdots\\ f_{m}\left(x_{1},\ldots,x_{n}\right)=0\end{cases}$
-is PR if for all finite partitions $\mathbb{N}=A_{1}\cup\ldots\cup A_{k}$ there are $i\leq k$ and $x_{1},\ldots,x_{n}\in A_{i}$ such that $\sigma\left(x_{1},\ldots,x_{n}\right)=0$.
-In [8], Schur proved one of the first and most important results in Arithmetic Ramsey Theory.
+we say that $\sigma\left(x_{1},\ldots,x_{n}\right)$ is PR if for every finite partition $\mathbb{N}=A_{1}\cup\ldots\cup A_{k}$, there are $i\leq k$ and $x_{1},\ldots,x_{n}\in A_{i}$ such that $\sigma\left(x_{1},\ldots,x_{n}\right)=\vec{0}$.
+In [8], Schur proved one of the first and most important results in rithmetic Ramsey theory.
 Theorem 1.2 (Additive Schur’s Theorem).
@@
 We just briefly recall a few definitions, mostly so that we can fix some notations.
-Sums and products of ultrafilters will be denoted by $\oplus$,$\odot$ , ⊙ respectively.
-Those notion are related to that of tensor product of ultrafilters in $\beta\mathbb{N}$, which is defined as follows:
+Sums and products of ultrafilters will be denoted by $\oplus$,$\odot$ , ⊙ respectively. They are related to tensor products of ultrafilters which, over $\mathbb{N}$, are defined as follows:
 Definition 2.1.
@@
 Moreover, we will need to talk about extensions of functions to $\beta\mathbb{N}$: for $f:\mathbb{N}^{n}\rightarrow\mathbb{N}$, we will denote by $\widehat{f}$ its unique continuous extension $\widehat{f}:\beta(\mathbb{N}^{n})\rightarrow\beta\mathbb{N}$.
-For example, when $S(x,y)=x+y$ is the sum, we have that $\hat{S}(p\odot q)=p\oplus q$ for all $p,q\in\beta\mathbb{N}$.
-As usual, to simplify notations we identify each natural number $n\in\mathbb{N}$ with the principal idempotent it generates, so that $\mathbb{N}\subseteq\beta\mathbb{N}$.
-As well-known, ultrafilters are often a fundamental tool to study PR problems. For configurations, the connection between ultrafilters and partition regularity reads as follows:
+As usual, to simplify notations we identify each natural number $n\in\mathbb{N}$ with the principal idempotent it generates, so that $\mathbb{N}\subseteq\beta\mathbb{N}$, and we convene that, for $a\in\mathbb{N},p\in\beta\mathbb{N}$, we will write $ap$ instead of $a\odot p$.
+As in Proposition 2.13 we will use ultrafilters over $\mathbb{R}^{+}_{d}$, we recall that $\beta\mathbb{R}^{+}_{d}$ is the Stone-Čech compactification of +:=$\mathbb{R}^{+}:=\{x\in\mathbb{R}\mid x>0\}$ endowed with the discrete topology.
+We adopt for $\beta\mathbb{R}^{+}_{d}$ the analogous notational conventions we are using for $\beta\mathbb{N}$.
+As well-known, ultrafilters are often a fundamental tool to study PR problems.
+For configurations, the connection between ultrafilters and partition regularity reads as follows (see e.g.
+[5, Theorem 5.7]):
 Fact 2.2.
@@
     $\exists p\in\beta\mathbb{N}\ \forall A\in p\ \exists x_{1},\ldots,x_{n}\in\mathbb{N}\ f_{1}\left(x_{1},\ldots,x_{n}\right),\ldots,f_{m}\left(x_{1},\ldots,x_{n}\right)\in A$.
-The above fact leads to the following definition:
+Analogously, for systems the following holds:
+Fact 2.3.
+Let $f_{1}\left(x_{1},\ldots,x_{n}\right),\ldots,f_{m}\left(x_{1},\ldots,x_{n}\right):\mathbb{N}^{n}\rightarrow\omega$, and let the system $\sigma\left(x_{1},\ldots,x_{n}\right)$ be defined by
+$\begin{cases}f_{1}\left(x_{1},\ldots,x_{n}\right)=0,\\ \ \ \ \ \ \ \ \ \ \vdots\\ f_{m}\left(x_{1},\ldots,x_{n}\right)=0.\end{cases}$
+The following are equivalent:
+  1. 1.
+    $\sigma\left(x_{1},\ldots,x_{n}\right)=\vec{0}$ is PR;
+  2. 2.
+    $\exists p\in\beta\mathbb{N}\ \forall A\in p\ \exists x_{1},\ldots,x_{n}\in A\ \sigma\left(x_{1},\ldots,x_{n}\right)=\vec{0}$.
+The above facts lead to the following definitions:
 Definition 2.4.
 Let $p\in\beta\mathbb{N}$ and let $f_{1}\left(x_{1},\ldots,x_{n}\right),\ldots,f_{m}\left(x_{1},\ldots,x_{n}\right):\mathbb{N}^{n}\rightarrow\mathbb{N}$.
-We say that $p$ witnesses the partition regularity of the configuration $\{f_{1}\left(x_{1},\ldots,x_{n}\right),\ldots,f_{m}\left(x_{1},\ldots,x_{n}\right)\}$ if for all $A\in p$ there exist $x_{1},\ldots,x_{n}\in A$ such that $f\left(x_{1},\ldots,x_{n}\right)=0$.
+We say that $p$ witnesses the partition regularity of the configuration $\{f_{1}\left(x_{1},\ldots,x_{n}\right),\ldots,f_{m}\left(x_{1},\ldots,x_{n}\right)\}$ if for all $A\in p$ there exist $x_{1},\ldots,x_{n}\in\mathbb{N}$ such that $f\left(x_{1},\ldots,x_{n}\right)\in A$.
 In this case, we write $p\models\{f_{1}\left(x_{1},\ldots,x_{n}\right),\ldots,f_{m}\left(x_{1},\ldots,x_{n}\right)\}$.
+Analogously, if the system $\sigma\left(x_{1},\ldots,x_{n}\right)$ is defined by
+$\begin{cases}f_{1}\left(x_{1},\ldots,x_{n}\right)=0,\\ \ \ \ \ \ \ \ \ \ \vdots\\ f_{m}\left(x_{1},\ldots,x_{n}\right)=0\end{cases}$
+we say that $p$ witnesses the partition regularity of $\sigma\left(x_{1},\ldots,x_{n}\right)=\vec{0}$ if $\forall A\in p\ \exists x_{1},\ldots,x_{n}\in A\ \sigma\left(x_{1},\ldots,x_{n}\right)=\vec{0}$.
+In this case, we write $p\models\sigma\left(x_{1},\dots,x_{n}\right)=\vec{0}$.
 In the next sections we will use a few specific results from other papers.
-For the readers easiness, we recall them here.
+For the readers’ easiness, we recall them here.
 The first states that when the PR of different systems of equations are witnessed by the same ultrafilter, these configurations can be mixed.
 Lemma 2.5.
-[1, Lemma 2.1] Let $p\in\beta\mathbb{N}$, and let $f\left(x_{1},\ldots,x_{n}\right),g\left(y_{1},\ldots,y_{m}\right)$ be two functions such that $p\models\left(f_{1}\left(x_{1},\ldots,x_{n}\right)=0\right)\wedge\left(g\left(y_{1},\ldots,y_{m}\right)=0\right)$.
-Then $p\models\left(f_{1}\left(x_{1},\ldots,x_{n}\right)=0\right)\wedge\left(g\left(y_{1},\ldots,y_{m}\right)=0\right)\wedge\left(x_{1}=y_{1}\right)$.
+[1, Lemma 2.1] Let $p\in\beta\mathbb{N}$, and let $f\left(x_{1},\ldots,x_{n}\right),g\left(y_{1},\ldots,y_{m}\right)$ be two functions such that $p\models\left(f\left(x_{1},\ldots,x_{n}\right)=0\right)\wedge\left(g\left(y_{1},\ldots,y_{m}\right)=0\right)$.
+Then $p\models\left(f\left(x_{1},\ldots,x_{n}\right)=0\right)\wedge\left(g\left(y_{1},\ldots,y_{m}\right)=0\right)\wedge\left(x_{1}=y_{1}\right)$.
 The second result that we will need is [4, Theorem 1.6], which states that multiplicatively piecewise syndetic sets contains some exponential structure.
@@
 [4, Theorem 1.6] Let $n\in\mathbb{N},n>1$ and let $A\subseteq\mathbb{N}$ be multiplicatively piecewise syndetic. Then there exists $x,y\in A$ such that $xn^{y}\in A$.
-The third result we need is [6, Theorem A], that states that certain ultrafilter equations admits no solution.
+The third and fourth result we need are [6, Theorem A] and (the restriction to $\mathbb{N}$ of) [5, Theorem 13.14] [^4] That we slightly rephrase to adapt to our notations., that states that certain ultrafilter equations admit no solution.
 Theorem 2.7.
 [6, Theorem A] Let $p\in\beta\mathbb{N}\setminus\mathbb{N}$, $a,b\in\mathbb{N}$ and $u,v\in\beta\mathbb{N}$. Suppose that $u\oplus ap=v\oplus bp$. Then $a=b$.
-Finally, we will the non-PR of combined sums and exponentiations stated in [7, Theorem 6].
+Theorem 2.8.
+[5, Theorem 13.14] Let $p,q,r,s\in\beta\mathbb{N}\setminus\mathbb{N}$. If |$|\{a\in\mathbb{N}\mid a\mathbb{Z}\in r\}|=|\mathbb{N}|$, then $q\oplus p\neq s\odot r$.
+Finally, we will use the non-PR of combined sums and exponentiations stated in [7, Theorem 6].
 Theorem 2.9.
@@
 We will provide some examples of nonprincipal ultrafilters $p,q$ satisfying $E_{1}(p,q)=E_{2}(q,p)$ after Proposition 2.13.
-To prove this proposition, it is helpful to use fractional powers of ultrafilters; to make proper sense of these powers, we prefer to state and prove Proposition 2.10 for ultrafilters in $\beta\mathbb{R}^{+}_{d}$, namely elements of the Stone-Čech compactification of +:=$\mathbb{R}^{+}:=\{x\in\mathbb{R}\mid x>0\}$ endowed with the discrete topology.
+To prove this proposition, it is helpful to use fractional powers of ultrafilters; to make proper sense of these powers, we prefer to state and prove Proposition 2.13 for ultrafilters in $\beta\mathbb{R}^{+}_{d}$.
 We extend the exponentiations to $\beta\mathbb{R}_{d}$ as follows: given $p,q\in\beta\mathbb{R}^{+}_{d}$, we let:
@@
 For $r\in\mathbb{R}^{+}$ and $p\in\beta\mathbb{R}^{+}_{d}$, we will keep the simplified notation $r^{p}$ to denote any of $E_{1}(r,p),E_{2}(p,r)$ (which are equal), and $p^{r}$ to denote any of $E_{1}(p,r),E_{2}(r,p)$ (which are also equal).
+We start with a preliminary lemma generalizing standard laws of exponents and logarithms.
+Lemma 2.12.
+Let $p\in\beta\mathbb{R}_{d}^{+}$ and let $a,b\in\mathbb{R}^{+}\setminus\{1\}$. The following facts hold:
+  1. 1.
+    $\widehat{\log_{a}}\left(b^{q}\right)=\log_{a}(b)\odot q$;
+  2. 2.
+    $\left(a^{b}\right)^{q}=a^{bq}$;
+  3. 3.
+    $\left(p^{a}\right)^{b}=p^{ab}$.
+Proof.
+We prove only $(1)$, as $(2)-(3)$ are similar. Let $A\subseteq\mathbb{R}^{+}$. Then
+$A\in\widehat{\log_{a}}\left(b^{q}\right)\Leftrightarrow\{x\in\mathbb{R}^{+}\mid\log_{a}(x)\in A\}\in b^{q}\Leftrightarrow$
+$\{z\in\mathbb{R}^{+}\mid b^{z}\in\{x\in\mathbb{R}^{+}\mid\log_{a}(x)\in A\}\}\in q\Leftrightarrow\{z\in\mathbb{R}^{+}\mid\log_{a}(b^{z})\in A\}\in q\Leftrightarrow$
+$\{z\in\mathbb{R}^{+}\mid z\cdot\log_{a}(b)\in A\}\in q\Leftrightarrow A\in\log_{a}(b)\odot q.\qed$
 Proposition 2.13.
@@
   2. 2.
-    $a^{p}=b^{q}\Leftrightarrow\left(\log_{a}(b)\right)p=q$;
+    $a^{p}=b^{q}\Leftrightarrow\left(\log_{a}(b)\right)\odot p=q$;
   3. 3.
@@
 The $E_{2}$-case is analogous.
-$(2)$ $\Rightarrow$)⇒ ) Consider the function $\log_{a}:\mathbb{R}^{+}\rightarrow\mathbb{R}$.
-A straightforward computation shows that $\widehat{\log_{a}}\left(a^{p}\right)=p$, hence our hypothesis entails that $p=\widehat{\log_{a}}\left(b^{q}\right)$.
-Now we observe that, for any $A\subseteq\mathbb{R}^{+}$,
-$A\in\widehat{\log_{a}}\left(b^{q}\right)\Leftrightarrow(\log_{a})^{-1}(A)\in\left(b^{q}\right)\Leftrightarrow\{x\in\mathbb{R}^{+}\mid\{y\in\mathbb{R}^{+}\mid x^{y}\in(\log_{a})^{-1}(A)\}\in q\}\in b\Leftrightarrow$
-$\{y\in\mathbb{R}^{+}\mid b^{y}\in(\log_{a})^{-1}(A)\}\in q\Leftrightarrow\{y\in\mathbb{R}^{+}\mid\log_{a}\left(b^{y}\right)\in A\}\in q\Leftrightarrow$
-$\{y\in\mathbb{R}^{+}\mid y\log_{a}(b)\in A\}\in q\Leftrightarrow A\in\left(\log_{a}(b)\right)q.$
-Hence $p=\left(\log_{a}(b)\right)q$, as desired.
-$\Leftarrow$)⇐ ) Just observe that $b^{q}=b^{\left(\log_{a}(b)\right)p}$, and $b^{\left(\log_{a}(b)\right)p}=E_{1}\left(b^{\left(\log_{a}(b)\right)},p\right)$ by $(1)$.
-We conclude as $b^{\left(\log_{a}(b)\right)}=a$.
-$(3)$ $\Rightarrow$)⇒ ) This time, consider the function $f(x):=\sqrt[a]{x}:\mathbb{R}^{+}\rightarrow\mathbb{R}$.
-Again, a straightforward computation shows that $\widehat{f}\left(p^{a}\right)=p$, hence $p=\widehat{f}\left(q^{b}\right)$.
-For any $A\subseteq\mathbb{R}^{+}$ we have that
-$A\in\widehat{f}\left(q^{b}\right)\Leftrightarrow(f)^{-1}(A)\in\left(q^{b}\right)\Leftrightarrow\{x\in\mathbb{R}^{+}\mid\{y\in\mathbb{R}^{+}\mid y^{x}\in f^{-1}(A)\}\in q\}\in b\Leftrightarrow$
-$\{y\in\mathbb{R}^{+}\mid y^{b}\in f^{-1}(A)\}\in q\Leftrightarrow\{x\in\mathbb{R}^{+}\mid f\left(x^{b}\right)\in A\}\in q\Leftrightarrow$
-$\{x\in\mathbb{R}^{+}\mid x^{\frac{b}{a}}\in A\}\in q\Leftrightarrow A\in q^{\frac{b}{a}}.$
-$\Leftarrow$)⇐ ) By $(1)$ we have that $q^{b}=E_{1}\left(p^{\frac{a}{b}},b\right)=E_{1}\left(p,\frac{a}{b}\cdot b\right)=p^{a}$.
+$(2)$ If $a^{p}=b^{q}$, by Lemma 2.12.(1) $p=\widehat{\log_{a}}\left(a^{p}\right)=\widehat{\log_{a}}\left(b^{q}\right)=\left(\log_{a}(b)\right)\odot q$; conversely, if $p=\left(\log_{a}(b)\right)\odot q$ then $a^{p}=a^{\left(\log_{a}(b)\right)\odot q}=\left(a^{\log_{a}(b)}\right)^{q}=b^{q}$ by Lemma 2.12.(2).
+$(3)$ If $p^{a}=q^{b}$, by Lemma 2.12.(3) $p^{\frac{a}{b}}=\left(p^{a}\right)^{\frac{1}{b}}=\left(q^{b}\right)^{\frac{1}{b}}=q$; conversely, again by Lemma 2.12.(3), if $p^{\frac{a}{b}}=q$ then $q^{b}=\left(p^{\frac{a}{b}}\right)^{b}=p^{\frac{ab}{b}}=p^{a}$.
 $(4)$ Let $A\subseteq\mathbb{R}^{+}$. Then
@@
 Corollary 2.14.
-Let $p,q,\beta\mathbb{N}\setminus{1}$, $a,b\in\mathbb{N}\setminus\{1\}$. Then
+Let $p,q\in\beta\mathbb{N}\setminus{1}$, $a,b\in\mathbb{N}\setminus\{1\}$. Then
   1. 1.
@@
 Example 2.15.
-Let $n\in\mathbb{N},n>1$ and $p,q\in\beta\mathbb{N}\setminus\{0\}$ be such that $p\odot q=q\odot p$. Then by Proposition 2.10.we have that
+Let $n\in\mathbb{N}\setminus\{1\}$ and $p,q\in\beta\mathbb{N}$ be such that $p\odot q=q\odot p$. Then by Proposition 2.13 we have that
 $E_{1}\left(n^{p},q\right)=n^{p\odot q}=n^{q\odot p}=E_{2}\left(q,n^{p}\right).$
-In particular, this shows the existence of infinitely many pairs of nonprincipal ultrafilters solving $E_{1}(p,q)=E_{2}(q,p)$ (just take, e.g., $p=2^{q}$).
+Let us notice that the equation $p\odot q=q\odot p$ can be solved by many pairs $p,q\in\beta\mathbb{N}\setminus\mathbb{N}$, for example by letting $p=q$, or by taking $p,q$ multiplicative idempotents with $p<q$.
+Hence, the equation $E_{1}(u,v)=E_{2}(u,v)$ has plenty of nontrivial solutions in $\beta\mathbb{N}\times\beta\mathbb{N}$.
 Proposition 2.13 might again suggest that many arithmetic properties of the exponentiations of natural numbers hold also for $E_{1},E_{2}$.
@@
 Proposition 2.16.
-Let $p\in\beta\mathbb{N}\setminus\{\mathbb{N}\}$, and let $a,b\in\mathbb{N}$. Then $p^{a}\odot p^{b}\neq p^{a+b}$.
+Let $p\in\beta\mathbb{N}\setminus\mathbb{N}$, and let $a,b\in\mathbb{N}$. Then $p^{a}\odot p^{b}\neq p^{a+b}$.
 Proof.
@@
 Let $F:\mathbb{N}\rightarrow\mathbb{N}$ be the function that maps a number $n$ to its largest prime divisor.
-Observe that, for all $p_{1},p_{2}\in\beta\mathbb{N}$, $\hat{\Omega}(E_{2}(p_{1},p_{2}))=p_{1}\odot\hat{\Omega}(p_{2})$ and $\hat{\Omega}(p_{1}\odot p_{2})=\hat{\Omega}(p_{1})\oplus\hat{\Omega}(p_{2})$.
+Observe that, for all $p_{1},p_{2}\in\beta\mathbb{N}$, $\widehat{F}(E_{2}(p_{1},p_{2}))=p_{1}\odot\widehat{F}(p_{2})$.
 We consider two cases.
-Case 1: $\hat{F}(p)=i\in\mathbb{N}$.
+Case 1: $\widehat{F}(p)=i\in\mathbb{N}$.
 In this case, consider the function $\Omega:\mathbb{N}\rightarrow\mathbb{N}$ that maps a number $n$ to the sum of the exponents in its prime factorization.
-Necessarily, $\hat{\Omega}(p)\notin\mathbb{N}$; in fact, if by contrast $\hat{\Omega}(p)=j\in\mathbb{N}$, we would have that $p\in\overline{S}$, where
+Necessarily, $\widehat{\Omega}(p)\notin\mathbb{N}$; in fact, if by contrast $\widehat{\Omega}(p)=j\in\mathbb{N}$, we would have that $p\in\overline{S}$, where
 $S=\{n\in\mathbb{N}\mid F(n)=i,\Omega(n)=j\},$
-which is a contradiction as $p$ is nonprincipal and $S$ is finite. Hence $\hat{\Omega}(p)\notin\mathbb{N}$. As $p^{a}\odot p^{b}=p^{a+b}$, we have that
-$\hat{\Omega}\left(p^{a}\odot p^{b}\right)=\hat{\Omega}\left(p^{a}\right)\oplus\hat{\Omega}\left(p^{b}\right)=a\hat{\Omega}(p)\oplus b\hat{\Omega},$
+which is a contradiction as $p$ is nonprincipal and $S$ is finite. Hence $\widehat{\Omega}(p)\notin\mathbb{N}$. As $p^{a}\odot p^{b}=p^{a+b}$, we have that
+$\widehat{\Omega}\left(p^{a}\odot p^{b}\right)=\widehat{\Omega}\left(p^{a}\right)\oplus\widehat{\Omega}\left(p^{b}\right)=a\widehat{\Omega}(p)\oplus b\widehat{\Omega}(p),$
 whilst
-$\hat{\Omega}\left(p^{a+b}\right)=(a+b)\hat{\Omega}(p).$
-We conclude as $a\hat{\Omega}(p)\oplus b\hat{\Omega}(p)\neq(a+b)\hat{\Omega}(p)$ by Theorem 2.6.
-Case 2: $\hat{F}(p)\notin\mathbb{N}$.
+$\widehat{\Omega}\left(p^{a+b}\right)=(a+b)\widehat{\Omega}(p).$
+We conclude as $a\widehat{\Omega}(p)\oplus b\widehat{\Omega}(p)\neq(a+b)\widehat{\Omega}(p)$ by Theorem 2.7.
+Case 2: $\widehat{F}(p)\notin\mathbb{N}$.
 Let $G:\mathbb{N}\rightarrow\mathbb{N}$ be the function $n\rightarrow\max\{m\in\mathbb{N}\mid F(n)^{m}\ \text{divides}\ n\}$, and let $H(n):=F(n)^{G(n)}$.
-As $\hat{F}(p)$ is nonprincipal, also $\hat{H}(p)$ is nonprincipal.
-And
-$\hat{H}\left(p^{a}\odot p^{b}\right)=b\hat{H}(p)\neq(a+b)\hat{H}(p)=\hat{H}\left(p^{a+b}\right),$
+As $\widehat{F}(p)$ is nonprincipal, also $\widehat{H}(p)$ is nonprincipal.
+Notice that:
+  • $\widehat{H}\left(p^{a}\odot p^{b}\right)=b\widehat{H}(p)$: in fact, for $x\in\mathbb{N}$, let $D_{x}=\{y\in\mathbb{N}\mid F(y)>F(x)\}$.
+If $x\in\mathbb{N}$ and $q\in\overline{D_{x}}$, then $\widehat{H}(x\odot q)=\widehat{H}(q)$.
+Therefore, if we let $\rho_{q}:\beta\mathbb{N}\rightarrow\beta\mathbb{N}$ be the function mapping $u\in\beta\mathbb{N}$ into $u\odot q$, we have that $\widehat{H}\circ\rho_{q}$ is constantly equal to $\widehat{H}(q)$ on $\mathbb{N}$; as both these functions are continuous and $\mathbb{N}$ is dense in $\beta\mathbb{N}$, $\widehat{H}(p\odot q)=\widehat{H}(q)$.
+Therefore $\widehat{H}\left(p^{a}\odot p^{b}\right)=\widehat{H}\left(p^{b}\right)=b\widehat{H}(p)$;
+  • $(a+b)\widehat{H}(p)=\widehat{H}\left(p^{a+b}\right)$, as for all $n\in\mathbb{N}$ $(a+b)H(n)=H\left(n^{a+b}\right)$.
+Therefore
+$\widehat{H}\left(p^{a}\odot p^{b}\right)=b\widehat{H}(p)\neq(a+b)\widehat{H}(p)=\widehat{H}\left(p^{a+b}\right),$
 again by Theorem 2.7. ∎
 Remark 2.17.
-The above results prove that the equalities $E_{1}(r,p)\odot E_{1}(r,q)=E_{1}(r,p\oplus q),E_{2}(p,r)\odot E_{2}(q,r)=E_{2}(p\oplus q,r)$ does not hold when $r$ is principal.
-They do not need to hold even when $r$ is nonprincipal.
-E.g., assume that $r$ is nonprincipal and $P=\{n\in\mathbb{N}\mid n\ \text{is a prime}\}\in r$.
+The above result proves that the equalities $E_{1}(p,a)\odot E_{1}(p,b)=E_{1}(p,a+b),E_{2}(a,p)\odot E_{2}(b,p)=E_{2}(a+b,p)$ does not hold when $a,b\in\mathbb{N},p\in\beta\mathbb{N}\setminus\mathbb{N}$.
+They do not need to hold even when $a,b\in\beta\mathbb{N}\setminus\mathbb{N}$.
+E.g., assume that $p,q,r$ are nonprincipal and $P=\{n\in\mathbb{N}\mid n\ \text{is a prime}\}\in p$.
 A direct computation shows that
-$I_{1}=\{n\in\mathbb{N}\mid\exists p_{1},p_{2}\in P,a,b\in\mathbb{N}\ \text{s.t.}\ p_{1}<a<p_{2}<b\ \text{and}\ p_{1}^{a}p_{2}^{b}=n\}\in E_{1}(r,p)\odot E_{1}(r,q),$
-$I_{2}=\{n\in\mathbb{N}\mid\exists p_{1},p_{2}\in P,a,b\in\mathbb{N}\ \text{s.t.}\ p_{1}>a>p_{2}>b\ \text{and}\ p_{1}^{a}p_{2}^{b}=n\}\in E_{2}(p,r)\odot E_{2}(q,r),$
+$I_{1}=\{n\in\mathbb{N}\mid\exists p_{1},p_{2}\in P,a,b\in\mathbb{N}\ \text{s.t.}\ p_{1}<a<p_{2}<b\ \text{and}\ p_{1}^{a}p_{2}^{b}=n\}\in E_{1}(p,q)\odot E_{1}(p,r),$
+$I_{2}=\{n\in\mathbb{N}\mid\exists p_{1},p_{2}\in P,a,b\in\mathbb{N}\ \text{s.t.}\ p_{1}>a>p_{2}>b\ \text{and}\ p_{1}^{a}p_{2}^{b}=n\}\in E_{2}(q,p)\odot E_{2}(r,p),$
 whilst
-$J=\{n\in\mathbb{N}\mid\exists p\in P,a\in\mathbb{N}\ \text{s.t.}\ p^{a}=n\}\in E_{1}(r,p\oplus q)\cap E_{2}(p\oplus q,r),$
-and we conclude that $E_{1}(r,p)\odot E_{1}(r,q)\neq E_{1}(r,p\oplus q)$ and $E_{2}(p,r)\odot E_{2}(q,r)\neq E_{2}(p\oplus q,r)$ as $I_{1}\cap J=I_{2}\cap J=\emptyset.$
-Their definition should make evident that $E_{1},E_{2}$ are ultrafilters operations that can be used to study the partition regularity of certain exponential configurations.
-Example 2.15.
-Let $p\in\beta\mathbb{N}\setminus\mathbb{N}$ and let $q=E_{1}(p,p)$. Let $A\in q$. Then
-$B_{0}=\{m\in\mathbb{N}\mid\{n\in\mathbb{N}\mid\{m^{n}\in A\}\in p\}\in p.$
-For $F_{1}\in\mathcal{P}_{f}(B_{0})$, let
-$B_{1}=\bigcap_{t\in F_{1}}\{n\in\mathbb{N}\mid t^{n}\in A\}\cap B_{0}\in p.$
-Proceeding inductively like above, it is possible to construct a sequence $\langle F_{i}\rangle_{i=1}^{\infty}$ of finite subsets of $B_{0}$ such that, for all $j<i$, $F_{j}^{F_{i}}=\{m^{n}\mid m\in F_{j},n\in F_{i}\}\in A$.
-For example, if $F_{i}=\{a_{i}\}$ for all $i$, we obtain an infinite subset $X=\{a_{i}\mid i\in\mathbb{N}\}$ of $B_{0}$ such that $\{a_{j}^{a_{i}}\mid a_{i},a_{j}\in X,j<i\}\subseteq A$.
-Arguing as above with $E_{2}$ instead of $E_{1}$, we could deduce an analogous results with the order relationship between $i$ and $j$ switched.
-The reader with a basic knowledge of ultrafilters will recognize that a trivial way to strengthen the above examples would be to take $p$ such that $E_{1}(p,p)=p$ or $E_{2}(p,p)=p$, namely some kind of exponential idempotent.
-We devote Section 2.3 to the study of the existence of such ultrafilters.
+$J=\{n\in\mathbb{N}\mid\exists p\in P,a\in\mathbb{N}\ \text{s.t.}\ p^{a}=n\}\in E_{1}(p,q\oplus r)\cap E_{2}(q\oplus r,p),$
+and we conclude that $E_{1}(p,q)\odot E_{1}(p,r)\neq E_{1}(p,q\oplus r)$ and $E_{2}(q,p)\odot E_{2}(r,p)\neq E_{2}(q\oplus r,p)$ as $I_{1}\cap J=I_{2}\cap J=\emptyset.$
 We now want to show that the operations $E_{1},E_{2}$ are related to the partition regularity of exponential triples, as well as of other exponential configurations.
-This should not come as a surprise: the main results in [2] can be restated in terminology introduced in Section 2.1 as follows:
+This should not come as a surprise: the main results in [2] can be restated in the terminology introduced in Section 2.1 as follows:
 Theorem 2.18 ([2, Theorem 2.1, restated]).
@@
 Theorem 2.19 ([2, Theorem 2.3, restated]).
-If $p$ is Van der Waerden’s witness then $2^{2^{p}\odot p}\models\{x,y,x^{y}\}$.
+If $p$ is a Van der Waerden’s witness then $2^{2^{p}\odot p}\models\{x,y,x^{y}\}$.
 As it is well-known (see e.g.
@@
   • by furtherly assuming that $p=p\odot p$, one gets that $q\models x^{\log_{2}(y)}=z$, hence e.g. $q\models x_{1}^{x_{2}\log_{2}x_{3}}=y_{1}\cdot y_{2}^{y_{3}}$;
-  • if $p$ is also combinatorially rich, then $p\models x=y+z^{2}$ (see e.g. [1, Theorem 2.11], hence $q\models x=yt^{\log_{2}(t)}$,
+  • if $p$ is also combinatorially rich [^6] See e.g. [1, Section 1.3] for a definition., then $p\models x=y+z^{2}$ (see e.g. [1, Theorem 2.11]), hence $q\models x=yt^{\log_{2}(t)}$,
 and so on, as we can always substitute a variable $x$ with a polynomial $P\left(y_{1},\ldots,y_{n}\right)$ whenever $q\models x=P\left(y_{1},\ldots,y_{n}\right)$.
@@
 Proposition 2.22.
-Let $q\in\beta\mathbb{N}\setminus\{0,1\}$. If $p$ is a Van der Waerden’s witness then $E_{1}\left(q,E_{1}(q,p)\odot p\right)\in\mathrm{ESW}$.
+Let $q\in\beta\mathbb{N}\setminus\{1\}$. If $p$ is a Van der Waerden’s witness then $E_{1}\left(q,E_{1}(q,p)\odot p\right)\in\mathrm{ESW}$.
 Proof.
@@
 Proposition 2.23.
-Let $p,q\in\beta\mathbb{N}\setminus\{0,1\}$. Assume that, for all $n>1$, $p\models\{x,y,xn^{y}\}$. Then $E_{1}\left(q,p\right)\in\mathrm{ESW}$.
+Let $p,q\in\beta\mathbb{N}\setminus\{1\}$. Assume that, for all $n>1$, $p\models\{x,y,xn^{y}\}$. Then $E_{1}\left(q,p\right)\in\mathrm{ESW}$.
 Proof.
@@
 Corollary 2.24.
-Let $q\in\overline{K(\beta\mathbb{N},\odot)}$. Then for all $p\in\beta\mathbb{N}\setminus\{0,1\}$ we have that $E_{1}\left(p,q\right)\in\mathrm{ESW}$.
+Let $q\in\overline{K(\beta\mathbb{N},\odot)}$. Then for all $p\in\beta\mathbb{N}\setminus\{1\}$ we have that $E_{1}\left(p,q\right)\in\mathrm{ESW}$.
 Remark 2.25.
-If $Sc_{+}=\{p\in\beta\mathbb{N}\mid\forall A\in p\ \exists a,b\in A\ a+b\in A\}$, Theorem 2.7 tells us that $Sc_{+}\cap\mathrm{ESW}=\emptyset$; in particular, if $E(\oplus)=\{p\in\beta\mathbb{N}\mid p\oplus p=p\}$ then
+If $\mathrm{ASW}=\{p\in\beta\mathbb{N}\mid\forall A\in p\ \exists a,b\in A\ a+b\in A\}$, Theorem 2.9 tells us that $\mathrm{ASW}\cap\mathrm{ESW}=\emptyset$; in particular, if $E(\oplus)=\{p\in\beta\mathbb{N}\mid p\oplus p=p\}$ then
 $\overline{K(\beta\mathbb{N},\odot)}\cap\mathrm{ESW}=\emptyset,\overline{E(\oplus)}\cap\mathrm{ESW}=\emptyset,$
-as $\overline{K(\beta\mathbb{N},\odot)}\cup\overline{E(\oplus)}\subseteq Sc_{+}$.
+as $\overline{K(\beta\mathbb{N},\odot)}\cup\overline{E(\oplus)}\subseteq\mathrm{ASW}$.
 By Theorem 2.23.(1), we get that if $q\in\overline{K(\beta\mathbb{N},\odot)}$ then $E_{1}\left(p,q\right)\notin\overline{K(\beta\mathbb{N},\odot)}$; in particular, for $p,q\in\overline{K(\beta\mathbb{N},\odot)}$, $E_{1}(p,q)\notin\overline{K(\beta\mathbb{N},\odot)}$: the smallest bilateral multiplicative ideal of $\beta\mathbb{N}$ is very much not closed under the $E_{1}$ exponentiation.
@@
 First of all, since $E_{1},E_{2}$ are not associative, we cannot approach the existence of idempotents problem via the usual method of applying Ellis’ Theorem (see [5, Theorem 2.5] for details).
-Moreover, since $1$ is the only principal $E_{1}-$idempotent, and the only $E_{2}$-idempotent as well, from now on we will use “idempotent” to actually mean “nonprincipal idempotent”.
+Moreover, since $1$ is the only principal $E_{1}-$idempotent, and the only principal $E_{2}$-idempotent as well, from now on we will use “idempotent” to actually mean “nonprincipal idempotent”.
 Our starting observation is the following:
 Proposition 2.27.
-Assume that there exist $p,q\in\beta\mathbb{N}\setminus\{0,1\}$ such that $p=p\odot E_{1}(q,p)$. Then $E_{1}(q,p)$ is $E_{1}$-idempotent.
-Analogously, if there exist $p,q\in\beta\mathbb{N}\setminus\{0,1\}$ such that $p=E_{2}(p,q)\odot p$ then $E_{2}(p,q)$ is $E_{2}$-idempotent.
+Assume that there exist $p,q\in\beta\mathbb{N}\setminus\{1\}$ such that $p=p\odot E_{1}(q,p)$. Then $E_{1}(q,p)$ is $E_{1}$-idempotent.
+Analogously, if there exist $p,q\in\beta\mathbb{N}\setminus\{1\}$ such that $p=E_{2}(p,q)\odot p$ then $E_{2}(p,q)$ is $E_{2}$-idempotent.
 Proof.
@@
 Proof.
-By [7, Lemma 23], there is a $4$-coloring $\mathbb{N}=A_{1}\cup\ldots\cup A_{4}$ of $\mathbb{N}$ such that there is no monochromatic pair $b,a^{b}$ with $\log_{2}a\leq b$ and $a,b\geq 1$.
+By [7, Lemma 23], there is a $4$-coloring $\mathbb{N}=A_{1}\cup\ldots\cup A_{4}$ of $\mathbb{N}$ such that there is no monochromatic pair $b,a^{b}$ with $\log_{2}a\leq b$ and $a,b>1$.
 Now assume that $E_{1}(p,q)=q$; by definition, this means that for all $A\subseteq\mathbb{N}$ we have that
@@
 $E_{2}$-idempotent ultrafilters do not exist.
-Unfortunately, we have not been able to prove the above conjecture; we can, however, show a partial result towards its solution.
+As expected, a positive answer to the above conjecture would have several interesting combinatorial consequences: for example, any set belonging to such an ultrafilter would be exponentially-IP, in the following sense.
+Theorem 2.31.
+Let $p\in\beta\mathbb{N}$ such that $p=E_{2}(p,p)$.
+For each $A\in p$, there is a sequence $\langle x_{n}\rangle_{n=1}^{\infty}$ in $\mathbb{N}$ such that $\{x_{n}\mid n\in\mathbb{N}\}\cup\{x_{n+1}^{y}\mid n\in\mathbb{N}\ \text{and}\ y\in FP(\langle x_{t}\rangle_{t=1}^{n})\}\subseteq A$.
+Proof.
+For $A\subseteq\mathbb{N}$ and $x\in\mathbb{N}$, let $D(A,x)=\{y\in\mathbb{N}\mid y^{x}\in A\}$; for each $A\in p$, $\{x\in\mathbb{N}\mid D(A,x)\in p\}\in p$ by $E_{2}$-idempotency of $p$.
+Let $A\in p$, let $A_{1}=A$, and let $B_{1}=\{x\in\mathbb{N}\mid D(A_{1},x)\in p\}$. Pick $x_{1}\in A_{1}\cap B_{1}$ and let $A_{2}=A_{1}\cap B_{1}\cap D(A_{1},x_{1})$.
+Now let $n\in\mathbb{N}\setminus\{1\}$ and assume that $A_{n}\in p$.
+Let $B_{n}=\{x\in\mathbb{N}\mid D(A_{n},x)\in p\}$, pick $x_{n}\in A_{n}\cap B_{n}$, and let $A_{n+1}=A_{n}\cap B_{n}\cap D(A_{n},x_{n})$.
+Having chosen $\langle x_{n}\rangle_{n=1}^{\infty}$, it is immediate that $\{x_{n}\mid n\in\mathbb{N}\}\subseteq A$.
+Now let $n\in\mathbb{N}$.
+We show by induction on $|F|$ that for $\emptyset\neq F\subseteq\{1,2,\dots,n\}$, if $m=\min F$ and $y=\prod_{t\in F}x_{t}$, then $x_{n+1}^{y}\in A_{m}$.
+If $F=\{m\}$, then $x_{n+1}\in A_{m+1}\subseteq D(A_{m},x_{m})$ so $x_{n+1}^{x_{m}}\in A_{m}$.
+Now assume that $|F|>1$, let $G=F\setminus\{m\}$, let $r=\min G$, and let $z=\prod_{t\in G}x_{t}$.
+Then $x_{n+1}^{z}\in A_{r}\subseteq A_{m+1}\subseteq D(A_{m},x_{m})$ so $x_{n+1}^{zx_{m}}=\left(x_{n+1}^{z}\right)^{x_{m}}\in A_{m}$.
+∎
+Unfortunately, we have not been able to prove Conjecture 2.30; we can, however, show a partial result towards its solution.
 The basic observation is the following: by Proposition 2.13, if $p$ is $E_{2}$-idempotent, we have that
@@
 Let $\Omega:\mathbb{N}\rightarrow\mathbb{N}$ be the function that maps $n$ to the sum of the exponents in its prime factorization.
-In particular, $\hat{\Omega}(p)=\hat{\Omega}(E_{2}(p,p))=p\odot\hat{\Omega}(p)$ and $\hat{\Omega}(p)=\hat{\Omega}(p\odot p)=\hat{\Omega}(p)\oplus\hat{\Omega}(p)$, therefore $\hat{\Omega}(p)$ is additively idempotent and $p\odot\hat{\Omega}(p)=\hat{\Omega}(p)\oplus\hat{\Omega}(p)$.
-If $\hat{\Omega}(p)\in\mathbb{N}$, then by additive idempotency we would get that $\hat{\Omega}(p)=0$, from which $p=1$, which is a contradiction as $p\in\beta\mathbb{N}\setminus\mathbb{N}$.
-So $\hat{\Omega}(p)$ is additively idempotent and nonprincipal, hence $n\mathbb{N}\in\hat{\Omega}(p)$ for all $n\in\mathbb{N}$.
-Thus we can apply [5, Theorem 13.14] to deduce that $p\odot\hat{\Omega}(p)\neq\hat{\Omega}(p)\oplus\hat{\Omega}(p)$, reaching the desired contradiction.
+In particular, $\widehat{\Omega}(p)=\widehat{\Omega}(E_{2}(p,p))=p\odot\widehat{\Omega}(p)$ and $\widehat{\Omega}(p)=\widehat{\Omega}(p\odot p)=\widehat{\Omega}(p)\oplus\widehat{\Omega}(p)$, therefore $\widehat{\Omega}(p)$ is additively idempotent and $p\odot\widehat{\Omega}(p)=\widehat{\Omega}(p)\oplus\widehat{\Omega}(p)$.
+If $\widehat{\Omega}(p)\in\mathbb{N}$, then by additive idempotency we would get that $\widehat{\Omega}(p)=0$, from which $p=1$, which is a contradiction as $p\in\beta\mathbb{N}\setminus\mathbb{N}$.
+So $\widehat{\Omega}(p)$ is additively idempotent and nonprincipal, hence $n\mathbb{N}\in\widehat{\Omega}(p)$ for all $n\in\mathbb{N}$.
+Thus we can apply Theorem 2.8 to deduce that $p\odot\widehat{\Omega}(p)\neq\widehat{\Omega}(p)\oplus\widehat{\Omega}(p)$, reaching the desired contradiction.
 ∎
@@
     $E_{2}(p,r)\odot E_{2}(q,r)=E_{2}(p\oplus q,r)$.
-For one ultrafilter, the following question, for which we conjecture a negative answer, seems relevant:
+One particular case of equality was discussed in Example 2.15, from which the following question arises:
 Question 3.3.
-Does it exist $p\in\beta\mathbb{N}\setminus\mathbb{N}$ such that $E_{1}(p,p)=E_{2}(p,p)$?
+For which $p,q\in\beta\mathbb{N}\setminus\mathbb{N}$ does the equality $E_{1}(p,q)=E_{2}(p,q)$ hold? Does it exist $p\in\beta\mathbb{N}\setminus\mathbb{N}$ such that $E_{1}(p,p)=E_{2}(p,p)$?
 Of course, as discussed in the paper, more relevant questions are the following:
@@
 What are the algebraical properties of $\mathrm{ESW}$?
+Acknowledgements: We thank the referee for his/her careful reading of the paper, which led to several improvements; in particular, we are thankful for the suggestion of isolating Lemma 2.12 to simplify the proofs of Proposition 2.13.(2)-(3), for proposing an explicit proof of the properties of $\widehat{H}$ used in Case 2 of Proposition 2.16, and for indicating Theorem 2.31 as a striking example of the eventual consequences of the existence of $E_{2}$-idempotents.